### v1

# Evaluating language models as risk scores

Abstract
Current question-answering benchmarks predominantly focus on accuracy in realizable prediction tasks. Conditioned on a question and answer-key, does the most likely token match the ground truth? Such benchmarks necessarily fail to evaluate language models’ ability to quantify outcome uncertainty. In this work, we focus on the use of language models as risk scores for unrealizable prediction tasks. We introduce folktexts, a software package to systematically generate risk scores using large language models, and evaluate them against benchmark prediction tasks. Specifically, the package derives natural language tasks from US Census data products, inspired by popular tabular data benchmarks. A flexible API allows for any task to be constructed out of 28 census features whose values are mapped to prompt-completion pairs. We demonstrate the utility of folktexts through a sweep of empirical insights on 16 recent large language models, inspecting risk scores, calibration curves, and diverse evaluation metrics. We find that zero-shot risk sores have high predictive signal while being widely miscalibrated: base models overestimate outcome uncertainty, while instruction-tuned models underestimate uncertainty and generate over-confident risk scores.
1 Introduction
Fueled by the success of large language models (LLMs), it is increasingly tempting for practitioners to use such models for risk assessment and decision making in consequential domains [1, 2, 3, 4]. Given the CV of a job applicant, for example, some might prompt a model, what are the chances that the employee will perform well on the job? The true answer is likely uncertain. Some applicants of the same features will do well, others won’t. A good statistical model should faithfully reflect such outcome uncertainty.
Calibration is perhaps the most basic kind of uncertainty quantification to ask for. A calibrated model, on average, reflects the true frequency of outcomes in a population. Calibrated models must therefore give at least some indication of uncertainty. Fundamental to statistical practice across the board, calibration has also been a central component in the debate around the ethics and fairness of consequential risk scoring in recent years [5, 6, 7, 8].
The evaluation of LLMs to date, however, has predominantly focused on accuracy metrics, often in realizable tasks where there is a unique correct label for each data point. Such benchmarks necessarily cannot speak to the use of a language model as risk scores. A model can have high utility in well-defined question-answering tasks while being wildly miscalibrated. Indeed, growing empirical evidence suggests that instruction-tuned language models tend to be over-confident in the probability that they assign to their preferred answer [9, 10].
Figure 1: Information flow from tabular data to risk scores, using a supervised classifier (left) or a language model (right). The folktexts package maps language models to the traditional machine learning workflow.
1.1 Our contributions
We contribute an open-source software package, called folktexts, [^1] Package and results available at: https://github.com/socialfoundations/folktexts that provides datasets and tools to evaluate statistical properties of LLMs on classification benchmarks. We show-case these functionalities with a sweep of new empirical insights on the risk scores produced by 16 recently proposed LLMs.
Classical metrics and tools for evaluating statistical properties of machine learning models rely on a standard type signature of numeric features $X$, labels $Y$, and risk scores $R$. As language models interface with the world strictly through natural text, they are incompatible with a rich repertoire of open-source libraries, benchmarks and evaluation tools. The folktexts package is designed to bridge this gap and offer a systematic way to translate between the natural language interface and the standard $(X,Y,R)$ statistical machine learning type signature. Figure 1 illustrates the workflow for producing risk scores using LLMs.
Inspired by the popular folktables package [11], folktexts derives natural-text machine learning datasets from US Census data products, specifically, the American Community Survey (ACS) [12]. Folktexts systematically constructs prompts from the survey responses of more than 3.2 million individuals, using the US Census codebook, and the ACS questionnaire as a reference. Risk scores are extracted from the model’s output token probabilities using a standard question-answering interface. The package offers five pre-specified question-answering benchmark tasks that are ready-to-use with any language model. These tasks can be customized through a flexible API where any census data attribute can be used either as a feature or as the target variable in a prediction task (see Table A5 for the initial set of 28 mapped attributes). Furthermore, evaluations can easily be performed over subgroups of the population to conduct algorithmic fairness audits.
Empirical insights.
We contribute a sweep of empirical findings based on our package. We evaluate 16 recently proposed LLMs, with sizes ranging from 2B parameters to 70B parameters. Our study demonstrates how inspecting risk scores of LLMS on underspecified prediction tasks reveals new insights that can not be deduced from inspecting accuracy alone. The main findings are summarized as follows:
  • Models’ output token probabilities have strong predictive signal, [^2] We measure predictive signal using the area under the receiver operating characteristic curve (AUC). but they are widely miscalibrated.
  • The failure modes of models are different: Base models consistently overestimate uncertainty and produce under-confident risk scores, while their instruction-tuned counterparts underestimate uncertainty and produce over-confident risk scores.
  • As a result instruction-tuning worsens calibration error on tasks with high outcome uncertainty, despite generally improving predictive signal and accuracy.
[…]
We hope our package facilitates future investigations into statistical properties of LLMs, not only as a way to faithfully reflect real-world uncertainties, but also as a required stepping stone to trustworthy model responses.
Outline.
In Section 2 we provide necessary background on risk scores and calibration in statistical machine learning. In Section 3 we extend this background to the application to language models, providing various design choices around prompting templates and ways to extract risk scores from language models. In Section 4 we evaluate 16 recent LLMs on 5 proposed benchmark tasks and summarize empirical findings.
1.2 Limitations
Predictive modeling and statistical risk scoring in consequential settings is a matter of active debate. Numerous scholars have cautioned us about the dangers of statistical risk scoring and, in particular, the potential of risk scores to harm marginalized and vulnerable populations [5, 13, 14, 15]. Our evaluation suite is intended to help in identifying one potential problem with language models for risk assessment, specifically, their inability to faithfully represent outcome uncertainty. However, our metrics are not intended to be sufficient criteria for the use of LLMs in consequential risk assessment applications. The fact that a model is calibrated says little about the potential impact it might have when used as a risk score. Numerous works in the algorithmic fairness literature, for example, have discussed the limitations of calibration as a fairness metric, see, e.g., [7, 8, 16]. There is also significant work on the limitations of statistical tools for predicting future outcomes and making decisions based on these predictions [17, 18, 19]. Calibration cannot and does not address these limitations.
1.3 Related work
The use of LLMs for decision-making has seen increasing interest as of late. Hegselmann et al. [20] show that a Bigscience T0 11B model [21] surpasses the predictive performance (AUC) of supervised learning baselines in the very-few-samples regime. The authors find that fine-tuning an LLM outperforms fitting a standard statistical model on a variety of small to medium-size tasks (from 768 to 48K samples) up to training set sizes in the 10s to 100s of samples.
Related work by Tamkin et al. [1] uses a similar tabular data set-up to study potential biases of the closed-source Claude 2.0 model. The authors generate hypothetical individual information for a variety of decision scenarios, and find that sensitive demographic information (such as age, race and gender) is actively used by the model for decision-making. Some attributes affect the decision positively (e.g., higher chance of approving a small business loan for minorities) while others affect it negatively (e.g., lower chance for older aged individuals).
It is unclear to which degree using synthetic or unrealistic data can affect the studies’ validity. Our proposed benchmarks enable similar studies to be conducted using real-world Census data. Moreover, related literature generally focuses on studying accurate class predictions by specific language models. We study over a dozen state-of-the-art language models, and discern which trends are actually consistent across all models, and which are constrained to a single model or model family. Further, we focus on a particularly neglected research avenue: analyzing LLM risk scores instead of discrete token answers, and whether they are accurate and calibrated.
Calibration.
Calibration is a widely studied concept in the literature on forecasting in statistics and econometrics with a venerable history [22, 23, 24, 25, 26, 27, 28]. Recent years have seen a surge in interest in calibration in the context of deep learning [29]. Calibration of LLMs has been studied on diverse question-answering benchmarks, ranging from sentiment classification, knowledge testing, and mathematical reasoning to multi-task benchmarks [e.g., 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. What differentiates our work is that we study calibration in naturally underspecified prediction tasks. This requires even the most accurate models to accurately reflect non-trivial probabilities over outcomes to be calibrated.
To systematically construct such prediction tasks we resort to survey data. Surveys have a long tradition in social science research as a tool for gathering statistical information about the characteristics and opinions of human populations [41]. Survey data comes with carefully curated questionnaires, as well as ground truth data. The value of this rich data source for model evaluation has not remained unrecognized. Surveys have recently gained attention to study bias and alignment of LLMs [42, 43, 44, 45, 46], and inspecting systematic biases in multiple-choice responses [47]. Instead of using surveys to get insights about a models natural inclinations, we use them to test model calibration with respect to a given population.
Beyond task-calibration, calibration at the word and token-level has been explored in the context of language generation [e.g., 48]. Others have focused on connections to hallucination [49], and expressing uncertainty in natural language [50, 51, 52]. Related to our work, Lin et al. [53] emphasize the inherent outcome uncertainty in language generation. Focusing on generation at the word or token level poses the challenge of measuring a high-dimensional probability distribution. Considering binary classification tasks has the practical advantage of circumventing this problem.
Calibration has also played a major role in the algorithmic fairness literature. Group-wise calibration has been proposed as a fairness criterion since the 1960s [54, 55]. In particular, it’s been a notion central to an active debate about the fairness of risk scores in consequential decision making [5, 6, 7, 8]. A recent line of work originating in algorithmic fairness studies multicalibration as a strengthening of calibration [56].
2 Preliminaries
This section provides necessary background on risk scores and the statistical evaluation of binary predictors. Throughout, we assume a joint distribution $\mathcal{P}$ given by a pair of random variables $(X,Y)$ where $X$ is a set of features and $Y$ is the outcome to be predicted. In the applications we study, the features $X$ typically form a text sequence and the outcome $Y$ is a discrete random variable that we would like to predict from the sequence. A parametric model $f_{\theta}(y|x)$ assigns a probability to each possible outcome $y$ given a feature vector $x$. The goal of a generative model is generally to approximate the conditional distribution $\mathcal{P}(y|x)$, where parameters are fit to a huge corpus of training data.
We will focus on binary prediction throughout this work. Let $Y\in\{0,1\}$ be a random variable indicating the outcome of an event the learner wishes to predict. We use the shorthand notation $f_{\theta}(x)$ to denote the model’s estimate of the probability that $Y=1$, given context $X=x$. Following standard terminology we will refer to $f_{\theta}:\mathcal{X}\rightarrow[0,1]$ as score function, and its output $f_{\theta}(x)$ as risk score. There are various dimensions along which to evaluate risk scores.
Calibration.
We say a score function is calibrated over a population $\mathcal{P}$ if and only if for all values $r\in[0,1]$ with $\mathbb{P}\{f_{\theta}(X)=r\}>0$, we have
$\mathbb{P}[Y=1\mid f_{\theta}(x)=r]=r.$ (1)
This condition asks that, over the set of all instances $x$ with score value $r$, an $r$ fraction of those instances must indeed have a positive label. Importantly, calibration is defined with respect to a population, it does not measure a model’s ability to discriminate between instances. A model that outputs the constant value $\mu=\operatorname{\mathbb{E}}[Y]$ on all instances is calibrated, by definition. In particular, we can always achieve calibration by setting $f_{\theta}(x)$ with the average value of $y$ among all instances $x^{\prime}$ in a given partition such that $f_{\theta}(x^{\prime})=f_{\theta}(x).$
We use Expected Calibration Error (ECE) as the primary metric to empirically evaluate calibration. It is defined as the expected absolute difference between a classifier’s confidence in its predictions and the accuracy on the same predictions. More formally, given $n$ triplets $(x_{i},y_{i},r_{i})$ where $r_{i}$ denotes the model’s score value for the corresponding data point $(x_{i},y_{i})$. The ECE is defined as
$\mathrm{ECE}\coloneq\frac{1}{n}\sum_{m}\left|\sum_{i\in B_{m}}y_{i}-\sum_{i\in B_{m}}r_{i}\right|,$ (2)
where data points are grouped into $M$ equally spaced bins $B_{m}$ according to their score values. We use $M=10$ in our evaluation, which is a commonly used value [57]. We also provide a quantile-based measure of ECE, as well as Brier score [22] to allow for a more complete picture. Furthermore, we use reliability diagrams [24, 28] to aid a visual interpretation of calibration. These diagrams plot expected sample accuracy as a function of confidence. Any deviation from a perfect diagonal represents miscalibration.
We can strengthen the calibration condition in (1) by requiring it in multiple subgroups of the population. Specifically, letting $G$ denote any discrete random variable, we can require the conditional calibration condition $\mathbb{P}[Y=1\mid G=g\,,f_{\theta}(x)=r]=r$ for every setting $g$ of the random variable $G.$ This is often used to define fairness.
Predictive signal.
The area under the receiver operating characteristic curve (AUC) is a rank-based measure of predictive model performance. It measures the probability that a randomly chosen positive observation ($Y=1$) will have a higher score than a randomly chosen negative observation ($Y=0$). A high AUC value in no way reflects accurate or calibrated probability estimates, it relates only to the signal-to-noise ratio in the risk scoring function [58]. Using the AUC metric allows us to neatly separate risk score calibration from their predictive signal, although both are crucial for accurate class predictions.
From risk scores to accuracy.
When solving classification problems it’s common practice to threshold risk scores to obtain a classifier. In our notation this corresponds to thresholding the risk scores $f_{\theta}(x)$:
$c(x)=\mathds{1}\{f_{\theta}(x)>\tau\}.$
The classifier $c$ that minimizes the misclassification error $\operatorname{\mathbb{E}}\mathds{1}\{c(X)\neq Y\}=\mathbb{P}\{c(X)\neq Y\}$ is given by $c^{*}(x)=\mathds{1}\{f^{*}(x)>0.5\}$, where $f^{*}$ is the Bayes optimal scoring function. In the following when we use accuracy we refer to the fraction of correct predictions after thresholding. If not specified otherwise we use $\tau=0.5$. This corresponds to an argmax operator applied to the class probabilities. It is important to note that a classifier can achieve perfect accuracy even when derived from a suboptimal scoring function. Thus, accuracy alone provides an incomplete picture of a model’s ability to express uncertainty.
3 Evaluating lanaguge models as risk scores
We are interested in the performance of LLMs as predictors in unrealizable binary prediction tasks. We prompt models with feature values $x$ to elicit risk scores $r$ and then evaluate these scores against ground truth labels $y$. Recall the illustration of the work flow in Figure 1.
3.1 Prediction tasks
We construct natural language prediction tasks from the American Community Survey (ACS) Public Use Microdata Sample (PUMS). [^3] https://www.census.gov/programs-surveys/acs/microdata.html The data contains survey responses of about 3.2 million anonymous individuals, carefully curated to offer statistical insights into the population of the United States. We refer to the data as Census data. The Census data contains demographic attributes, as well as information related to income, employment, health, transportation, and housing. Prediction tasks are defined by selecting a subset of attributes to define the features and one attribute to be the label. We threshold continuous target variables and bin multi-class predictions to obtain a binary classification task. To enable straightforward comparison with existing tabular benchmarks, we consider natural text analogues to the tasks in the popular folktables benchmark package [11]. Appendix B describes each task in further detail.
Natural uncertainty in risk scoring.
Prediction tasks on human populations typically come with natural outcome uncertainty, meaning that the target label is not uniquely determined by the input features. The fewer features are provided the higher the uncertainty in the outcome. We take this to our advantage to systematically evaluate model calibration. Namely, to be calibrated, a risk score has to reflect both model uncertainty and uncertainty inherent to the prediction task. In fact, the optimal predictor would often output low-confidence answers. In contrast, prevailing question-answering benchmarks always require high confidence for an accurate model to be calibrated. Underspecified, non-realizable prediction tasks allow us to circumvent such potential confounding in calibration evaluations.
3.2 Extracting risk scores
To extract risk scores from LLMs we map each inference problem to a natural text prompt. For a given datapoint, we specify the classification task in the prompt and extract class probabilities from the model’s next token probabilities, similar to traditional question-answering interfaces. The prompt consists of three components:
  • Instantiating population: We first instantiate the population $\mathcal{P}$ in the prompt context. This corresponds to the population represented by our reference data. We use third-person prompting: “The following data describes a survey respondent. The survey was conducted among US
    residents in 2018. Please answer the question based on the information provided.
    ” This step is typically not needed in supervised classification tasks, because the training data implicitly defines the population. However, for LLMs this is particularly important for evaluating calibration outside the realizable setting, as risk score can not in general simultaneously be calibrated to different populations. Skipping this step could provide insights related to alignment [42, 43] rather than calibration.
[…]
Using the constructed prompts, we query the models and extract scores from the next token probabilities for the choice labels A,B as $r_{i}=\mathbb{P}(\texttt{A})/(\mathbb{P}(\texttt{A})+\mathbb{P}(\texttt{B}))$, following the methodology of standard question-answering benchmarks [36, 42]. As LLMs are known to have ordering biases in multiple-choice question-answering [45, 62], we evaluate responses on all choice orderings and average the resulting scores. We refer to Xiong et al. [33] for an overview on alternative design choices on how to elicit confidence scores from LLMs.
3.3 The folktexts package
The folktexts package is designed to offer a flexible interface between tabular prediction tasks and natural language question-answering tasks in order to extract risk scores from LLMs. The package makes available the ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, ACSTravelTime prediction tasks [11] as natural language benchmarks, together with various functionalities to customize the task definitions and subsample the reference data. The set of attributes available to define the features and label can be found in the ACS PUMS data dictionary.444 https://www.census.gov/programs-surveys/acs/microdata/documentation.html
In addition to providing a reproducible way to extract risk scores from LLMs, folktexts also offers pre-implemented evaluation metrics to benchmark and compare the calibration and accuracy of LLMs. To offer researchers a quick overview over model performance the package also offers functionalities to plot calibration curves for visual inspection of results. The package is easy to use within a python notebook, or from the command line. Further details on the design choices and architecture are available in Appendix C. The next section demonstrates the utility of the package.
4 Empirical findings
We use folktexts to evaluate several recently released models together with their instruction-tuned counterparts: the Llama 3 models [63], including the 8B and the 70B versions, the Mistral 7B [64], the Mixtral 8x7B and 8x22B variants [65], the Yi 34B [66], and the Gemma [67] 2B and 7B variants. All LLM results are obtained via zero-shot prompting. For comparison, results are also shown for a logistic regression (LR) model, and a gradient boosted decision trees model (XGBoost). The XGBoost model [68] is generally regarded as the state-of-the-art in tabular data tasks [69, 70].
In this section we focus on the ACSIncome prediction task, which is the default folktexts benchmarking task. It consists in predicting whether a person’s income is above or below \$50K from 10 demographic features, and closely emulates the popular UCI Adult prediction task [71]. The evaluation test set consists of 160K randomly selected samples from the 2018 Census data. A separate set of 1.5M samples is used to train the supervised LR and XGBoost models, LLMs are used as zero-shot classifiers without fine-tuning. Experiments were ran on a cluster with Nvidia-A100-80GB GPUs, consuming an approximate total of 500 GPU hours. Appendix A presents additional results and plots for the experiments analyzed in this section, including the study of alternate prediction tasks.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.27 0.27 0.86 0.69 0.78 0.42
Llama 3 70B 0.20 0.20 0.86 0.70 0.77 0.14
Llama 3 8B (it) 0.32 0.30 0.85 0.62 0.78 0.37
Llama 3 8B 0.25 0.26 0.81 0.38 0.69 0.05
Mixtral 8x22B (it) 0.21 0.22 0.85 0.76 0.75 0.48
Mixtral 8x22B 0.17 0.19 0.85 0.68 0.77 0.10
Mixtral 8x7B (it) 0.16 0.18 0.86 0.78 0.78 0.42
Mixtral 8x7B 0.17 0.21 0.83 0.65 0.74 0.06
Mistral 7B (it) 0.21 0.22 0.83 0.77 0.77 0.42
Mistral 7B 0.20 0.23 0.80 0.73 0.74 0.04
Yi 34B (chat) 0.19 0.19 0.86 0.72 0.74 0.34
Yi 34B 0.24 0.22 0.85 0.62 0.74 0.20
Gemma 7B (it) 0.61 0.59 0.84 0.37 0.75 0.06
Gemma 7B 0.24 0.27 0.76 0.37 0.68 0.04
Gemma 2B (it) 0.63 0.63 0.73 0.37 0.63 0.00
Gemma 2B 0.14 0.25 0.62 0.45 0.63 0.02
LR 0.03 0.18 0.79 0.74 - -
XGBoost 0.00 0.13 0.90 0.82 - -
Table 1: Zero-shot LLM results on the ACSIncome benchmark task, together with results for LR and XGBoost baselines fitted on 1.5M samples. Column “Acc. $\tau=t^{*}$” shows accuracy results after fitting a model binarization threshold on $n=100$ train samples, while “Acc. $\tau=0.5$” shows standard accuracy using the highest likelihood answer. Table cells are colored using a continuous color map between the worst ( in orange) and best ( in cyan) results of each column, where only the top 10% and bottom 10% of the scale is colored. Models generally achieve high predictive signal (high AUC) but poor calibration (high ECE).
4.1 Benchmark results
First, we perform a comprehensive evaluation of risk scores output by LLMs along multiple metrics. Results are summarized in Table 1.
We first compare LLMs to the supervised learning baselines. We observe that a majority of LLMs (all of size 8B or larger) outperform the linear model in terms of area under the AUC curve. However, LLMs are clearly far from matching the supervised baselines with respect to calibration: all language models achieve high calibration error (ECE), while baselines achieve near-perfect calibration. Due to this high miscalibration, models struggle to translate scores with high predictive signal into high accuracy. We recall that AUC is agnostic to calibration, while accuracy on the maximum likelihood answer is not. All of the Gemma models have worse than random accuracy, and most LLMs fail to surpass even the linear baseline on this metric. Only the instruction-tuned Mistral models (7B, 8x7B, and 8x22B) outperform the linear model in terms of accuracy (see accuracy with $\tau=0.5$).
To show-case potential accuracy gains from improved calibration, we add the “Acc. $\tau=t^{*}$” column, where the risk-score binarization threshold $\tau\in[0,1]$ is fitted on a small set of $n=100$ random training samples. As expected, LLMs with poor calibration but high AUC see steep accuracy improvements after re-calibrating the threshold; e.g., Gemma 7b (it) jumps from 0.37 to 0.75 accuracy. Inspecting accuracy alone would fail to reveal the predictive signal of the risk scores.
Finally, we focus on comparing base models to their instruction-tuned counterparts, marked with ‘(it)’ or ‘(chat)’. A striking trend is visible across the board: instruction-tuning generally worsens calibration. At the same time, we see small but consistent improvements in AUC after instruction-tuning. Appendix A.1 presents results on the four additional prediction tasks. The same trend of instruction-tuning leading to worse calibration and higher AUC is broadly replicated. However, performance across different tasks is somewhat inconsistent: LLMs generally have stronger predictive signal than a supervised linear model on the income and travel-time prediction tasks, but consistently underperform the linear baseline on the address change (mobility) task.
Figure 2: Calibration curves for base (left) and instruction-tuned (right) models of the Llama and Mistral families. Curves are computed using $10$ quantile-based score bins.
4.2 Score distribution
To get additional insights into the difference between base models and their instruction-tuned counterparts, we inspect their risk scores more closely.
Figure 2 shows calibration curves of base and instruction-tuned models of the Llama and Mistral families. Both base and instruction-tuned models have poor score calibration, but failure modes are entirely different: Base models output under-confident scores, while instruction-tuned models output over-confident scores. To quantify this result, we introduce a measure of risk score confidence bias similar to the ECE metric:
$\mathrm{R}_{\mathrm{bias}}\coloneq\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left[\mathrm{Conf}(B_{m})-\mathrm{Acc}(B_{m})\right],$
where $M$ is the number of score bins, $B_{m}$ is the set of samples in score bin $m$, $\mathrm{Acc}(.)$ is the accuracy on a given set of samples, and $\mathrm{Conf}(.)$ is the confidence on a given set of samples measured as the mean risk score for the highest likelihood class. Figure 3 shows the risk score confidence bias results. The two miscalibration modes are evident: confidence bias is higher for instruction-tuned models and lower (or even negative) for base models. That is, instruction-tuned models are generally over-confident in their predictions, outputting higher scores than their accuracy would warrant, while no such trend is visible for base models. Section 4.3 further analyzes subgroup calibration and its implications to algorithmic fairness.
Figure 4 shows the risk score distribution for a variety of model pairs. The score distributions for base and instruct model variants are immediately distinguishable: base models consistently produce low-variance distributions centered around 0.5, while instruction-tuned variants often output scores near 0 or 1. Even for the two versions of Llama 3 70B, which achieve exactly the same AUC, the score distributions are widely different. On this model pair, no predictive performance is gained by instruction tuning, but answers of the instruction-tuned model have significantly higher (over-)confidence. In fact, while the base Llama 3 70B has an average of 0.07 under-confidence bias, the instruct variant produces risk scores on average 0.21 over-confident (see Figure 3). The same trend is visible on all model pairs, with Yi 34B showing the smallest difference between base and instruct variants. Note that the instruct Gemma models degenerate into predicting only positive outcomes with high confidence score, hence why they have the lowest accuracy, worst Brier score, and worst ECE.
Interestingly, in contrast to knowledge testing where high confidence is typically associated with better models, our evaluation reveals that even if the true outcome has high entropy, instruction-tuning polarizes score distributions.
Figure 3: Risk score confidence bias for all LLMs on the ACSIncome task. Negative values indicate under-confident risk scores (overestimating uncertainty), while positive values indicate over-confident risk scores (underestimating uncertainty). For all but the Yi model pair, the instruct model shows higher over-confidence.
Figure 4: Risk score distribution for base and instruction-tuned model pairs on the ACSIncome task. After instruction-tuning, models exhibit high confidence, but worse calibration in general. The XGBoost scores showcase a perfectly calibrated distribution ($\text{ECE}\approx 0.00$). Consult Table 1 for details.
Figure 5: Shift of score distribution with increasing evidence on the ACSIncome task, for the Mixtral 8x7B model (which achieved the best Brier score). Features are described in Table A5. Score distribution gets more discriminative as more evidence is added, successfully increasing scores’ predictive signal (AUC).
4.3 Subgroup calibration
Next, we evaluate risk score calibration across different subpopulations. Figure 6 shows calibration curves for two sets of models on the ACSIncome task, evaluated on three subgroups specified by the instantiation of the race attribute in the data. We pick the three race categories with the largest representation in the ACS data. Appendix A.2 presents further discussion on subgroup miscalibration and its potential causes and implications. Note that a positive prediction of $Y=1$ is arguably the advantageous outcome, as it corresponds to the high-income category (“Earns above \$50,000 per year”).
[…]
Figure 6: Calibration curves for the Yi 34B and the Mixtral 8x22B models, across different race subpopulations on the ACSIncome task. Left: Base models. Right: Instruction-tuned models. The instruction-tuned models shown are worse offenders in terms of subgroup-specific score bias. The Yi 34B models (top) overestimate income, but the Black population sees consistently lower overestimates. The Mixtral 8x22B models (bottom) both over- and underestimate income depending on the score value, and output lower scores for the Black population.
4.4 Varying degree of uncertainty
Finally, we consider the dependence of risk scores on the available evidence. For this study we use the Mixtral 8x7B model, which achieves the best (lowest) Brier score among evaluated models (reflecting both high accuracy and high calibration). Figure 5 shows score distributions with increasing evidence: starting with only 2 features, and iteratively adding 2 features at a time. This sequence demonstrates how unrealizable, underspecified prediction tasks differ from realizable prediction tasks. Predicting income based on an individual’s place of birth (POBP) and race (RAC1P) is naturally not possible to a high degree of accuracy, forcing any calibrated model to output lower-confidence risk scores. Indeed, both base and instruct variants correctly output lower confidence scores for the smaller feature sets (left plots) when compared with the larger feature sets (right plots). Appendix A.3 goes further in-depth on how score distributions change with varying data uncertainty.
In addition to providing insights into risk scores, varying individual features in the prediction task also provides insights into LLM feature importance. Appendix A.4 presents LLM feature importance results and discusses the main differences to traditional supervised learning model.
5 Conclusion
We introduced folktexts, a software package that provides datasets and tools to evaluate risk scores produced by language models. We believe that folktexts strengthens the evaluation ecosystem in a direction that was previously underserved. Unlike most existing LLM benchmarks, the datasets we introduced have inherent outcome uncertainty, making them useful as a basis for systematic evaluation of uncertainty quantification in LLMs.
At present, our empirical findings show that LLM risk scores generally have strong predictive signal, but are wildly miscalibrated. Such models may be good for knowledge testing, but lack adequate indicators of uncertainty. Additionally, we show that most instruction-tuned models have worse calibration but slightly higher accuracy and AUC than their base-model counterparts. Crucially, we demonstrate how inspecting calibration curves and score distributions is necessary to understand differences across models, such as the differences between base and instruct model variants, which are obfuscated when using single calibration metrics.
Going forward, folktexts features a modular, easily extensible architecture. The package therefore makes it easy to evaluate different models, prompting strategies, and metrics. We define reproducible reference standards for evaluating LLMs, but encourage researchers to use what they believe is best for different applications. We envision that future package extensions will include other uncertainty quantification methods, such as confidence intervals, or conformal prediction methods.
Acknowledgments
We thank Florian Dorner, Mila Gorecki, and Ricardo Dominguez-Olmedo for invaluable feedback on an earlier version of this paper. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting André F. Cruz.
Supplementary Materials A Additional results
This appendix section shows additional results and corresponding plots to support the insights presented in Section 4. Section A.1 shows results on four extra benchmark tasks made available with folktexts. Section A.2 extends the discussion on subgroup calibration and algorithmic fairness on the ACSIncome task. Section A.3 goes further in-depth on how to use folktexts to control data uncertainty in the benchmark prediction tasks. Finally, Section A.4 presents and discusses results on feature importance for LLM predictions.
A.1 Results on additional benchmark tasks
The main body of the paper focuses on results on the ACSIncome prediction task. This task is arguably the most popular for benchmarking on tabular data, as it closely mirrors the older but widely used UCI Adult dataset [11, 71]. The folktexts package makes available natural-language versions of four additional tabular data tasks: ACSEmployment, ACSMobility, ACSTravelTime, and ACSPublicCoverage.
Tables A1A4 show results for the ACSEmployment, ACSMobility, ACSTravelTime, and ACSPublicCoverage tasks, respectively. Trends discussed in the main body of the paper are broadly confirmed. Models’ moderate predictive performance is accompanied by substantial miscalibration of their risk scores. Additionally, base models output low-variance high-uncertainty score distributions, while instruction-tuned models output high-variance low-uncertainty score distributions. There is clear predictive signal for large models across all tasks (e.g., see the AUC of Llama 3 70B it, or Mixtral 8x22B it). However, the extent to which models’ scores are predictable varies substantially across tasks. Most models surpass the AUC of the linear baseline (LR) on the ACSTravelTime task, as well as the main ACSIncome task; but consistently lag behind the linear baseline on the ACSMobility task. On the ACSEmployment and ACSPublicCoverage tasks, the best performing models manage to match the linear baseline AUC.
These findings pose into question one of the main advantages of using LLMs for risk scoring: the fact that no labeled data is required. Given the inconsistency of model performance, some small amount of testing data may always be needed to assert reliability of results.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.17 0.19 0.85 0.73 0.78 0.29
Llama 3 70B 0.25 0.26 0.82 0.51 0.74 0.12
Llama 3 8B (it) 0.07 0.19 0.79 0.74 0.74 0.22
Llama 3 8B 0.34 0.34 0.75 0.45 0.68 0.06
Mixtral 8x22B (it) 0.38 0.39 0.60 0.51 0.60 0.26
Mixtral 8x22B 0.21 0.24 0.86 0.52 0.77 0.08
Mixtral 8x7B (it) 0.22 0.23 0.83 0.73 0.75 0.44
Mixtral 8x7B 0.29 0.31 0.81 0.45 0.74 0.06
Mistral 7B (it) 0.35 0.36 0.72 0.63 0.66 0.49
Mistral 7B 0.26 0.30 0.76 0.45 0.69 0.04
Yi 34B (chat) 0.14 0.21 0.79 0.69 0.73 0.20
Yi 34B 0.08 0.23 0.70 0.62 0.63 0.11
Gemma 7B (it) 0.36 0.38 0.59 0.58 0.59 0.44
Gemma 7B 0.16 0.25 0.65 0.48 0.64 0.13
Gemma 2B (it) 0.38 0.41 0.42 0.46 0.48 0.23
Gemma 2B 0.01 0.24 0.63 0.54 0.61 0.03
LR 0.02 0.15 0.86 0.78 - -
XGBoost 0.00 0.12 0.91 0.83 - -
Table A1: Zero-shot LLM results on the ACSEmployment benchmark task, together with supervised learning baselines fitted on 2.9M samples. See caption of Table 1 for more details.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.20 0.25 0.57 0.61 0.48 0.18
Llama 3 70B 0.22 0.24 0.55 0.67 0.53 0.04
Llama 3 8B (it) 0.15 0.22 0.56 0.70 0.53 0.17
Llama 3 8B 0.10 0.20 0.55 0.73 0.63 0.04
Mixtral 8x22B (it) 0.41 0.40 0.51 0.40 0.50 0.24
Mixtral 8x22B 0.11 0.20 0.55 0.73 0.63 0.02
Mixtral 8x7B (it) 0.26 0.26 0.58 0.73 0.60 0.02
Mixtral 8x7B 0.14 0.21 0.57 0.73 0.52 0.03
Mistral 7B (it) 0.26 0.26 0.57 0.73 0.54 0.02
Mistral 7B 0.20 0.23 0.53 0.73 0.33 0.01
Yi 34B (chat) 0.09 0.20 0.56 0.72 0.59 0.15
Yi 34B 0.07 0.20 0.57 0.73 0.63 0.08
Gemma 7B (it) 0.25 0.26 0.58 0.73 0.67 0.11
Gemma 7B 0.41 0.37 0.50 0.39 0.41 0.10
Gemma 2B (it) 0.73 0.73 0.52 0.27 0.40 0.00
Gemma 2B 0.25 0.26 0.51 0.39 0.45 0.02
LR 0.02 0.19 0.61 0.74 - -
XGBoost 0.00 0.16 0.74 0.76 - -
Table A2: Zero-shot LLM results on the ACSMobility benchmark task, together with supervised learning baselines fitted on 0.6M samples. See caption of Table 1 for more details.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.15 0.24 0.70 0.60 0.61 0.12
Llama 3 70B 0.09 0.24 0.67 0.55 0.61 0.05
Llama 3 8B (it) 0.19 0.28 0.60 0.57 0.57 0.11
Llama 3 8B 0.08 0.25 0.53 0.56 0.54 0.03
Mixtral 8x22B (it) 0.32 0.33 0.66 0.59 0.61 0.15
Mixtral 8x22B 0.20 0.28 0.63 0.44 0.54 0.04
Mixtral 8x7B (it) 0.45 0.45 0.66 0.52 0.55 0.29
Mixtral 8x7B 0.28 0.32 0.60 0.44 0.56 0.03
Mistral 7B (it) 0.41 0.42 0.59 0.57 0.53 0.19
Mistral 7B 0.05 0.25 0.57 0.56 0.55 0.02
Yi 34B (chat) 0.35 0.36 0.65 0.56 0.52 0.02
Yi 34B 0.08 0.24 0.62 0.56 0.52 0.06
Gemma 7B (it) 0.42 0.43 0.53 0.56 0.50 0.02
Gemma 7B 0.04 0.24 0.61 0.58 0.57 0.02
Gemma 2B (it) 0.34 0.36 0.49 0.56 0.54 0.03
Gemma 2B 0.09 0.26 0.48 0.44 0.45 0.01
LR 0.04 0.24 0.58 0.56 - -
XGBoost 0.02 0.19 0.77 0.70 - -
Table A3: Zero-shot LLM results on the ACSTravelTime benchmark task, together with supervised learning baselines fitted on 1.3M samples. See caption of Table 1 for more details.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.16 0.21 0.69 0.75 0.49 0.20
Llama 3 70B 0.18 0.22 0.67 0.63 0.50 0.07
Llama 3 8B (it) 0.11 0.21 0.59 0.71 0.71 0.19
Llama 3 8B 0.41 0.38 0.55 0.30 0.50 0.02
Mixtral 8x22B (it) 0.24 0.25 0.70 0.72 0.67 0.17
Mixtral 8x22B 0.32 0.30 0.59 0.30 0.56 0.05
Mixtral 8x7B (it) 0.20 0.23 0.70 0.74 0.72 0.35
Mixtral 8x7B 0.41 0.37 0.57 0.30 0.38 0.03
Mistral 7B (it) 0.30 0.30 0.61 0.70 0.68 0.01
Mistral 7B 0.29 0.30 0.45 0.30 0.70 0.02
Yi 34B (chat) 0.06 0.19 0.67 0.74 0.68 0.13
Yi 34B 0.04 0.21 0.59 0.70 0.56 0.07
Gemma 7B (it) 0.30 0.34 0.46 0.50 0.32 0.32
Gemma 7B 0.15 0.23 0.49 0.49 0.37 0.10
Gemma 2B (it) 0.70 0.70 0.54 0.30 0.33 0.00
Gemma 2B 0.26 0.28 0.54 0.30 0.62 0.01
LR 0.03 0.19 0.70 0.73 - -
XGBoost 0.00 0.14 0.84 0.80 - -
Table A4: Zero-shot LLM results on the ACSPublicCoverage benchmark task, together with supervised learning baselines fitted on 1.0M samples. See caption of Table 1 for more details.
A.2 Additional subgroup calibration results
This section contains additional subgroup calibration plots on the ACSIncome task. Figure 6 shows the 2 worst offenders: Yi 34B and Mixtral 8x22B. Figures A1 and A2 show the remaining models. The same trend is visible but to a lesser extent: Instruction-tuned models show a bias towards lower scores for Black individuals, even after controlling for label prevalence.
We can quantify the positive score bias by evaluating the signed calibration error (SCE):
$\mathrm{SCE}\coloneq\frac{1}{n}\sum_{m=1}^{M}\sum_{i\in B_{m}}(r_{i}-y_{i}),$ (3)
where $M$ is the number of score buckets, $B_{m}$ is the set of sample indices belonging to bucket $m$, $r_{i}=f_{\theta}(x_{i})$ is the risk score given to sample $x_{i}$ and $y_{i}$ is its label. This metric does not evaluate overall calibration, as a value of 0 does not indicate a calibrated classifier. Instead, negative/positive values indicate a bias towards lower/higher risk scores, respectively. If higher scores are related to positive real-world outcomes (e.g., when predicting income for a loan application), then a bias towards lower scores on samples of specific protected subgroups would likely lead to unfair real-world outcomes. Conversely, if a negative class prediction is associated with a positive outcome (e.g., when predicting risk of recidivism), then a bias towards lower scores would be beneficial for the affected group.
Figure A3 shows the difference between the signed calibration error (SCE) on different group pairings, on the ACSIncome task. Positive predictions ($Y=1$) correspond to the advantaged high-income class. Negative differences, $\Delta_{SCE}<0$, indicates advantaged scores for Black individuals, while positive differences, $\Delta_{SCE}>0$, indicate disadvantaged scores for Black individuals. Interestingly, while subgroup calibration curves appear similar for base models and disadvantage Black individuals for instruction-tuned models (see Figures 6A1, and A2), this is not entirely reflected on the SCE metric. Indeed, a clear-cut split is visible: base models benefit the score of Black individuals, and instruct models disadvantage the score of Black individuals. This finding could be partially explained by the fact that base models produce score distributions with low variance, and instruct models produce high-variance polarized outcomes. Specifically, two conclusions can be drawn from the score distributions produced by base models: (1) models under-estimate the score of high-income earners (which are disproportionately Asian), and (2) models over-estimate the score of low-income earners (which are disproportionately Black). This could lead base models to over-estimate the earnings of the Black population disproportionately to other groups. The opposite is true for instruction-tuned models: high-income earners see their score over-estimated, which benefits groups with a higher prevalence of high earners.
This in no way serves as an exhaustive analysis of risk score fairness on LLMs, as it is bound to be highly task dependent and language model dependent. We simply surface the fact that, on this high-income prediction task, risk scores are not group-calibrated [8] and could lead to unfair outcomes. Crucially, even though race has the lowest mean feature importance among all features (see Appendix A.4), we report and explain how different trends in risk score distributions can effectively lead to unfair outcomes.
Figure A1: Calibration curves for the Mistral family of models, across different race sub-populations on the ACSIncome task. Left: Base models. Right: Instruction-tuned models.
Figure A2: Calibration curves for the Gemma family of models, across different race sub-populations on the ACSIncome task. Left: Base models. Right: Instruction-tuned models. Note that these are the smallest models and generally perform the worst on all tasks. The plotted calibration curves are in agreement with the lackluster performance, as they show extreme score distributions that are unlike those of other LLMs. In fact, the instruction-tuned Gemma models consistently output only positive predictions, hindering their analysis.
Figure A3: Racial group bias in risk score calibration error. When comparing score bias of group $A$ with score bias of group $B$, $\Delta_{\mathrm{SCE}}=\mathrm{SCE}_{A}-\mathrm{SCE}_{B}$, positive values indicate an undue score advantage of group $A$, and negative values an undue score advantage of group $B$. Left: Difference between White and Black groups, $\mathrm{SCE}_{White}-\mathrm{SCE}_{Black}$. Right: Difference between Asian and Black groups, $\mathrm{SCE}_{Asian}-\mathrm{SCE}_{Black}$. Note that the Gemma models were omitted, as their instruct versions degenerate into strictly predicting only positive outcomes (see Figure A2). Consequently, the two instruct Gemma models are the only exceptions to the trend shown in these plots.
A.3 Varying uncertainty
A simple API call in our package allows for selecting different subsets of attributes to include as features when using the LLM as a predictor. Figure A4 inspects the effect of increasing the feature on models’ calibration and predictive power. Each dot along the line represents an increasing feature set used for LLM predictions, added in order of mean feature importance on all models. Appendix B details all features used in the ACSIncome task. We refer the reader to Ding et al. [11] and the ACS codebook [^5] See the ACS PUMS data dictionary for the full list of available variables:
https://www.census.gov/programs-surveys/acs/microdata/documentation.html
for an in-depth description of each categorical value a feature can take.
As expected, predictive signal increases when using larger feature sets, but not in a straightforward way: some features actually worsen performance. This contrasts with standard supervised learning algorithms, where all features would either be neutral or contribute positively to a model’s predictive power when tested on data that is i.i.d. with the training data. On the other hand, there is no clear trend for calibration: for Mistral models, it seems that calibration actually worsens for larger feature sets, while for Llama models calibration is approximately stable across all points.
This experiment show-cases one unique way of using LLMs with survey prediction tasks: while supervised learning models would have to be retrained every time a different feature set is used, LLMs can freely change the evidence they use to make a prediction. If a model were to exhibit properties of a joint distribution with the ability to marginalize over hidden features, then calibration with respect to evidence $\mathcal{X}$ implies that it is also calibrated with respect to restricted evidence $\mathcal{X}^{\prime}\subset\mathcal{X}$. To what extent a model satisfies such properties is an interesting question for future work; we hope our package proves useful as an investigative tool.
Figure A4: Evaluation of calibration (ECE) and predictive performance (AUC) on Llama and Mistral models, with an increasing number of features provided on the ACSIncome task. For each dot along the line we add two features, up to all 10 features being used in the point marked with a star. Top row: base models. Bottom row: instruction-tuned models. Models can successfully use each extra feature to increase predictive signal. Calibration trends worse the more features are added for base models, while instruct models show no clear trend.
A.4 Feature importance
In this section, we present feature importance results for different LLMs on the ACSIncome prediction task. The importance value of feature $j$ is computed as the drop in AUC after permuting all values of feature $j$ across the dataset. That is, each sample $x$, sees its value for feature $j$ randomly permuted with another sample. This is a common feature importance implementation [72], as it does not rely on any internal characteristics of the model.
Figure A5 shows feature importance values for the largest language models studied (above 40B active parameters). Results for the XGBoost model are also shown in green. Note that XGBoost achieves the best result on every single metric in Table 1. While for supervised models, a given categorical value is nothing more than a 1 or 0, LLMs have the potential to surface the real-world meaning of such values, benefiting from the rich embedding representations of each category. As such, we’d expect to see LLMs assigning higher importance to categorical features. Indeed, Llama 3 models assign considerably higher importance to the occupation feature (OCCP), which is a numerically encoded categorical feature with over 500 different possible values. Conversely, the XGBoost model assigns considerably higher importance than LLMs to ‘work-hours per week’ (WKHP) and ‘age’ (AGEP), both integer-encoded features. Lastly, feature importance results indicate that the studied LLMs do not explicitly use sensitive categories such as age (AGEP), sex (SEX), or race (RAC1P) for risk score estimation.
Interestingly, feature importance is similar for base and instruct variants of the same model. This contrasts with the score distribution and calibration curve results, where all base models followed a similar trend, distinct from their instruction-tuned versions.
Figure A5: Feature importance among the largest language models tested, plus results for the XGBoost baseline. Feature importance values are calculated as the loss in AUC when the values of a given column are randomly permuted [72].
Supplementary Materials B Details on provided benchmark tasks
The folktexts package defines natural-text mappings for a variety of columns in the ACS PUMS data files. Table A5 lists and describes each implemented column-to-text mapping. Any combination of column-to-text objects can be used to create a prediction task from ACS data, both as features and as the prediction target. To enable straightforward comparison with existing benchmarks, we mimic the feature set and population filters used by the prediction tasks available in the popular folktables benchmark package [11]. Specifically, we put forth natural-text variants of the ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, and ACSTravelTime tasks. These prediction tasks define a restricted set of columns from the ACS PUMS data files to be used as input features for machine learning models, as well as a binarized target column. As such, we extend the use of these ACS prediction tasks to benchmark language models, enabling direct comparison with a wide-ranging set of literature works. Although any ACS survey year could be used for benchmarking, we define the standard set of benchmark tasks as those using data from the 2018 1-year-horizon person-level survey (following Ding et al. [11]). Notwithstanding, we welcome the addition of new column-to-text mappings by new users of the package, both for ACS data and for new datasets. The following paragraphs detail each pre-implemented prediction task.
ACSIncome
The goal of the ACSIncome task is to predict whether a person’s yearly income is above \$50,000, given by the PINCP column. The ACS columns used as features are: AGEP, COW, SCHL, MAR, OCCP, POBP, RELP, WKHP, SEX, and RAC1P. The sub-population over which the task is conducted is employed US residents with age greater than 16 years. The ACSIncome prediction task was put-forth as the successor to the popular UCI Adult dataset [71], used extensively in the algorithmic fairness literature. This task is the default task when running the folktexts benchmark.
ACSPublicCoverage
The goal of the ACSPublicCoverage task is to predict whether an individual is covered by public health insurance, given by the PUBCOV column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, CIT, MIG, MIL, ANC, NATIVITY, DEAR, DEYE, DREM, PINCP, ESR, ST, FER, and RAC1P. The sub-population over which the task is conducted is US residents with age below 65 years old, and with personal income below \$30,000.
ACSMobility
The goal of the ACSMobility task is to predict whether an individual has changed their home address in the last year, given by the MIG column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, CIT, MIL, ANC, NATIVITY, RELP, DEAR, DEYE, DREM, RAC1P, COW, ESR, WKHP, JWMNP, and PINCP. The sub-population over which the task is conducted is US residents with age between 18 and 35.
ACSEmployment
The goal of the ACSEmployment is to predict whether an individual is employed, given by the ESR column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, MIG, CIT, MIL, ANC, NATIVITY, RELP, DEAR, DEYE, DREM, and RAC1P. The sub-population over which the task is conducted is US residents with age between 16 and 90.
ACSTravelTime
The goal of the ACSTravelTime task is to predict whether a person’s commute time to work is greater than 20 minutes, given by the JWMNP column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, MIG, RELP, RAC1P, ST, CIT, OCCP, JWTR, and POVPIP. The sub-population over which the task is conducted is employed US residents with age greater than 16 years.
Supplementary Materials C folktexts package usage
The folktexts package is made available to the public via its open-source code repository1 and as a standalone package to be installed via the Python Package Index (PyPI). The main user-facing classes are Benchmark, BenchmarkConfig, LLMClassifier, TaskMetadata, ColumnToText, and Dataset. The responsibilities of each class are ascribed as follows
  • The Benchmark class is responsible for running a benchmark task, which consists in obtaining risk scores from a given LLM on a given dataset, and evaluating those predictions on a variety of benchmark metrics.
  • The BenchmarkConfig class details all configurations of a benchmark (see Figure A6 for available options).
  • The LLMClassifier class is comprised of a transformers model, a tokenizer, and a task; and is responsible for producing risk scores given some tabular rows for the provided task.
  • The TaskMetadata class is responsible for defining a set of feature columns and target column, together with holding the corresponding column-to-text objects to map an entire tabular row to its natural-text representation. The benchmark ACS tasks instantiate a subclass named ACSTaskMetadata.
  • The ColumnToText class is responsible for producing meaningful natural-text representations of each possible value of a numeric or categorical column.
  • The Dataset class is responsible for holding tabular data and enabling reproducible manipulation of that data, such as splitting in train/test/validation, or filtering for a specified sub-population. The data used for the benchmark ACS tasks is provided by a subclass named ACSDataset.
Additionally, a command-line interface is provided to ease usability: The benchmark ACS tasks can be ran using the run_acs_benchmark executable. Figure A6 details each available flag. Further infromation and example notebooks can be found on github at: https://github.com/socialfoundations/folktexts.
usage: run_acs_benchmark [-h] model MODEL task-name TASK_NAME results-dir RESULTS_DIR data-dir DATA_DIR [–few-shot FEW_SHOT] [–batch-size BATCH_SIZE] [–context-size CONTEXT_SIZE] [–fit-threshold FIT_THRESHOLD] [–subsampling SUBSAMPLING] [–seed SEED] [–dont-correct-order-bias] [–chat-prompt] [–direct-risk-prompting] [–reuse-few-shot-examples] [–use-feature-subset [USE_FEATURE_SUBSET …]]
[–use-population-filter [USE_POPULATION_FILTER …]] [–logger-level {DEBUG,INFO,WARNING,ERROR,CRITICAL}]
Run a benchmark folktexts task.
options:
-h, help show this help message and exit
model MODEL [str] Model name or path to model saved on disk
task-name TASK_NAME
[str] Name of the ACS task to run the experiment on
results-dir RESULTS_DIR
[str] Directory under which this experiments results will be saved
data-dir DATA_DIR [str] Root folder to find datasets on
few-shot FEW_SHOT [int] Use few-shot prompting with the given number of shots
batch-size BATCH_SIZE
[int] The batch size to use for inference
context-size CONTEXT_SIZE
[int] The maximum context size when prompting the LLM
fit-threshold FIT_THRESHOLD
[int] Whether to fit the prediction threshold, and on how many samples
subsampling SUBSAMPLING
[float] Which fraction of the dataset to use (if omitted will use all data)
seed SEED [int] Random seed to set for reproducibility
dont-correct-order-bias
[bool] Whether to avoid correcting ordering bias, by default will correct it
chat-prompt [bool] Whether to use chat-based prompting (for instruct models)
direct-risk-prompting
[bool] Whether to directly prompt for risk-estimates instead of multiple-choice Q&A
reuse-few-shot-examples
[bool] Whether to reuse the same samples for few-shot prompting (or sample new ones every time)
use-feature-subset [USE_FEATURE_SUBSET …]
[str] Optional subset of features to use for prediction
use-population-filter [USE_POPULATION_FILTER …]
[str] Optional population filter for this benchmark; must follow the format column_name=value to filter the dataset by a specific value.
logger-level {DEBUG,INFO,WARNING,ERROR,CRITICAL}
[str] The logging level to use for the experiment
Figure A6: Documentation for using folktexts package through the command-line interface. An executable named run_acs_benchmark is made available to run the standard ACS benchmark tasks.
Column Description Example
AGEP Age The individual’s age is: 42 years old.
COW Class of worker The individual’s current employment status is: Working for a non-profit organization.
SCHL Educational attainment The individual’s highest grade completed is: 12th grade.
MAR Marital status The individual’s marital status is: Married.
OCCP Occupation The individual’s occupation is: Human Resources Manager.
POBP Place of birth The individual’s place of birth is: New Zealand.
RELP Relationship The individual’s relationship to the reference survey respondent in the household is: Brother or sister.
WKHP Work-hours per week The individual’s usual number of hours worked per week is: 40 hours.
SEX Sex The individual’s sex is: Female.
RAC1P Race The individual’s race is: Black or African American.
PINCP Total yearly income The individual’s total yearly income is: \$75,000.
CIT Citizenship status The individual’s citizenship status is: Naturalized US citizen.
DIS Disability status The individual has a disability.
ESP Employment status of parents The individual is living with two parents: both parents in labor force.
MIG Mobility (lived here 1 year ago) The individual lived in the same house 1 year ago.
MIL Military service The individual was on active duty in the past, but not currently.
PUBCOV Public health coverage The individual is covered by public health insurance.
ANC Ancestry The individual has single ancestry.
NATIVITY Nativity The individual is foreign born.
DEAR Hearing The individual has hearing difficulty.
DEYE Vision The individual does not have vision difficulty.
DREM Cognition The individual does not have cognitive difficulties.
ESR Employment status #2 The individual is not in the labor force.
ST State The individual lives in California.
FER Parenthood (1 year) The individual gave birth to a child within the past 12 months.
JWMNP Commute time The individual takes 45 minutes travelling to work every day.
JWTR Means of transport The individual’s means of transport to work is a bicycle.
POVPIP Income-to-poverty ratio The individual’s income to poverty ratio is 150%.
Table A5: Description of all column-to-text mappings implemented for ACS features. The variable part of each example is shown in typeset grey font. Details on each possible categorical value for each feature are available in the ACS PUMS data dictionary.5