### v3

# Evaluating language models as risk scores

Abstract
Current question-answering benchmarks predominantly focus on accuracy in realizable prediction tasks. Conditioned on a question and answer-key, does the most likely token match the ground truth? Such benchmarks necessarily fail to evaluate LLMs’ ability to quantify ground-truth outcome uncertainty. In this work, we focus on the use of LLMs as risk scores for unrealizable prediction tasks. We introduce folktexts, a software package to systematically generate risk scores using LLMs, and evaluate them against US Census data products. A flexible API enables the use of different prompting schemes, local or web-hosted models, and diverse census columns that can be used to compose custom prediction tasks. We evaluate 17 recent LLMs across five proposed benchmark tasks. We find that zero-shot risk scores produced by multiple-choice question-answering have high predictive signal but are widely miscalibrated. Base models consistently overestimate outcome uncertainty, while instruction-tuned models underestimate uncertainty and produce over-confident risk scores. In fact, instruction-tuning polarizes answer distribution regardless of true underlying data uncertainty. This reveals a general inability of instruction-tuned LLMs to express data uncertainty using multiple-choice answers. A separate experiment using verbalized chat-style risk queries yields substantially improved calibration across instruction-tuned models. These differences in ability to quantify data uncertainty cannot be revealed in realizable settings, and highlight a blind-spot in the current evaluation ecosystem that folktexts covers.
1 Introduction
Fueled by the success of large language models (LLMs), it is increasingly tempting for practitioners to use such models for risk assessment and decision making in consequential domains [1, 2, 3, 4]. Given the CV of a job applicant, for example, some might prompt a model, what are the chances that the employee will perform well on the job? The true answer is likely uncertain. Some applicants of the same features will do well, others won’t. A good statistical model should faithfully reflect such outcome uncertainty.
Calibration is perhaps the most basic kind of uncertainty quantification to ask for. A calibrated model, on average, reflects the true frequency of outcomes in a population. Calibrated models must therefore give at least some indication of uncertainty. Fundamental to statistical practice across the board, calibration has also been a central component in the debate around the ethics and fairness of consequential risk scoring in recent years [5, 6, 7, 8].
The evaluation of LLMs to date, however, has predominantly focused on accuracy metrics, often in realizable tasks where there is a unique correct label for each data point. Such benchmarks necessarily cannot speak to the use of language models as risk score estimators. A model can have high utility in well-defined question-answering tasks while being wildly miscalibrated. In fact, while accuracy corresponds to knowledge over the expected answer, proper uncertainty quantification corresponds to knowledge over the variance over answers.
Figure 1: Information flow from tabular data to risk scores, using a supervised classifier (left) or a language model (right). The folktexts package maps language models to the traditional machine learning workflow.
1.1 Our contributions
We contribute an open-source software package, called folktexts, [^1] Package and results available at: https://github.com/socialfoundations/folktexts that provides datasets and tools to evaluate statistical properties of LLMs as risk scorers. We show-case its functionalities with a sweep of new empirical insights on the risk scores produced by 17 recently proposed LLMs.
The folktexts package offers a systematic way to translate between the natural language interface and the standard machine learning type signature. It translates prediction tasks defined by numeric features $X$ and labels $Y$ into natural text prompts and extracts risk scores $R$ from LLMs. This opens up a rich repertoire of open-source libraries, benchmarks and evaluation tools to study their statistical properties. Figure 1 illustrates the workflow for producing risk scores using LLMs.
For benchmarking risk scores, we need ground truth samples from a known probability distribution. Inspired by the popular folktables package [9], folktexts builds on US Census data products, specifically, the American Community Survey (ACS) [10], collecting information about more than 3.2 million individuals representative of the US population. Folktexts systematically constructs prediction tasks and prompts from the individual survey responses using the US Census codebook, and the ACS questionnaire as a reference. Risk scores are extracted from the language model’s output token probabilities using a standard question-answering interface. The package offers five pre-specified question-answering benchmark tasks that are ready-to-use with any language model. A flexible API allows for a variety of natural language tasks to be constructed out of 28 census features whose values are mapped to prompt-completion pairs (features detailed in Table A5). Furthermore, evaluations can easily be performed over subgroups of the population to conduct algorithmic fairness audits.
Empirical insights.
We contribute a sweep of empirical findings based on our package. We evaluate 17 recently proposed LLMs, with sizes ranging from 2B parameters to 141B parameters. Our study demonstrates how inspecting risk scores of LLMs on underspecified prediction tasks reveals new insights that cannot be deduced from inspecting accuracy alone. The main findings are summarized as follows:
  • Models’ output token probabilities have strong predictive signal [^2] We measure predictive signal using the area under the receiver operating characteristic curve (AUC) [11]. but are widely miscalibrated.
  • The failure modes of models are different: Multiple-choice answer probabilities generated by base models consistently overestimate outcome uncertainty, while instruction-tuned models underestimate uncertainty and produce over-confident risk scores.
  • Using a verbalized chat-style prompt results in materially different answer distributions, with significantly improved calibration for instruction-tuned models, accompanied by a small but consistent decrease in predictive power.
  • Instruction-tuning generally worsens calibration of multiple-choice answers, but improves calibration of verbalized numeric answers.
We hope our package facilitates future investigations into statistical properties of LLMs, not only as a way to faithfully reflect real-world uncertainties, but also as a required stepping stone to trustworthy model responses.
Outline.
In Section 2 we provide necessary background on risk scores and calibration in statistical machine learning. In Section 3 we extend this background to the application to language models, providing various design choices around prompting templates and ways to extract risk scores from language models. In Section 4 we evaluate 17 recent LLMs on 5 proposed benchmark tasks and summarize empirical findings.
1.2 Limitations
Predictive modeling and statistical risk scoring in consequential settings is a matter of active debate. Numerous scholars have cautioned us about the dangers of statistical risk scoring and, in particular, the potential of risk scores to harm marginalized and vulnerable populations [5, 12, 13, 14]. Our evaluation suite is intended to help in identifying one potential problem with language models for risk assessment, specifically, their inability to faithfully represent outcome uncertainty. However, our metrics are not intended to be sufficient criteria for the use of LLMs in consequential risk assessment applications. The fact that a model is calibrated says little about the potential impact it might have when used as a risk score. Numerous works in the algorithmic fairness literature, for example, have discussed the limitations of calibration as a fairness metric, see, e.g., [7, 8, 15]. There is also significant work on the limitations of statistical tools for predicting future outcomes and making decisions based on these predictions [16, 17, 18]. Calibration cannot and does not address these limitations.
1.3 Related work
The use of LLMs for decision-making has seen increasing interest as of late. Hegselmann et al. [19] show that a Bigscience T0 11B model [20] surpasses the predictive performance (AUC) of supervised learning baselines in the very-few-samples regime. The authors find that fine-tuning an LLM outperforms fitting a standard statistical model on a variety of tasks up to training set sizes in the 10s to 100s of samples. Related work by Slack and Singh [21] shows that providing task-specific expert knowledge via instructions in context can lead to significant improvements in model predictive power. Tamkin et al. [1] generate hypothetical individual information for a variety of decision scenarios, and analyze how language models’ outputs change when provided different demographic data. Some attributes are found to positively affect the model’s decision (e.g., higher chance of approving a small business loan for minorities) while others affect it negatively (e.g., lower chance for older aged individuals).
A separate research thread considers how to leverage LLMs to model human population statistics. Argyle et al. [22] evaluate whether GPT-3 can faithfully reproduce political party preferences for different US subpopulations, and conclude that model outputs can accurately reflect a variety of correlations between demographics and political preferences. Other works use a similar methodology to model the distribution of human opinions on different domains [23, 24, 25]. Aher et al. [26], Dillion et al. [27] use LLMs to reproduce popular psychology experiments on human moral judgments, and confirm there’s good alignment between human answers and LLM outputs. Literature on modelling human population statistics generally focuses on using LLMs to obtain accurate survey completions. That is, given demographic information on an individual, what was their response to a specific survey question? This methodology often ignores the fact that individuals described by the same set of demographic features will realistically give different answers. An accurate model would not only provide the highest likelihood answer, but also a measure of uncertainty corresponding to the expected variability within a sub-population. Our work tackles this arguably neglected research avenue: Analyzing LLM risk scores instead of discrete token answers, and whether they are accurate and calibrated to human populations.
Calibration.
Calibration is a widely studied concept in the literature on forecasting in statistics and econometrics with a venerable history [28, 29, 30, 31, 32]. Recent years have seen a surge in interest in calibration in the context of deep learning [33]. Calibration of LLMs has been studied on diverse question-answering benchmarks, ranging from sentiment classification, knowledge testing, and mathematical reasoning to multi-task benchmarks [e.g., 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. What differentiates our work is that we study calibration in naturally underspecified prediction tasks. This requires even the most accurate models to accurately reflect non-trivial probabilities over outcomes to be calibrated.
To systematically construct such prediction tasks we resort to survey data. Surveys have a long tradition in social science research as a tool for gathering statistical information about the characteristics and opinions of human populations [44]. Survey data comes with carefully curated questionnaires, as well as ground truth data. The value of this rich data source for model evaluation has not remained unrecognized. Surveys have recently gained attention to study bias and alignment of LLMs [22, 45, 46, 47, 48], and inspecting systematic biases in multiple-choice responses [49]. Instead of using surveys to get insights about a models natural inclinations, we use them to test model calibration with respect to a given population.
Beyond task-calibration, calibration at the word and token-level has been explored in the context of language generation [e.g., 50]. Others have focused on connections to hallucination [51], and expressing uncertainty in natural language [52, 53, 54]. Related to our work, Lin et al. [55] emphasize the inherent outcome uncertainty in language generation. Focusing on generation at the word or token level poses the challenge of measuring a high-dimensional probability distribution. Considering binary classification tasks has the practical advantage of circumventing this problem.
Calibration has also played a major role in the algorithmic fairness literature. Group-wise calibration has been proposed as a fairness criterion since the 1960s [56, 57]. In particular, it’s been a notion central to an active debate about the fairness of risk scores in consequential decision making [5, 6, 7, 8]. A recent line of work originating in algorithmic fairness studies multicalibration as a strengthening of calibration [58].
2 Preliminaries
This section provides necessary background on risk scores and the statistical evaluation of binary predictors. Throughout, we assume a joint distribution $\mathcal{P}$ given by a pair of random variables $(X,Y)$ where $X$ is a set of features and $Y$ is the outcome to be predicted. In the applications we study, the features $X$ typically form a text sequence and the outcome $Y$ is a discrete random variable that we would like to predict from the sequence. A parametric model $f_{\theta}(y|x)$ assigns a probability to each possible outcome $y$ given a feature vector $x$. The goal of a generative model is generally to approximate the conditional distribution $\mathcal{P}(y|x)$, where parameters are fit to a huge corpus of training data.
We will focus on binary prediction throughout this work. Let $Y\in\{0,1\}$ be a random variable indicating the outcome of an event the learner wishes to predict. We use the shorthand notation $f_{\theta}(x)$ to denote the model’s estimate of the probability that $Y=1$, given context $X=x$. Following standard terminology we will refer to $f_{\theta}:\mathcal{X}\rightarrow[0,1]$ as score function, and its output $f_{\theta}(x)$ as risk score. There are various dimensions along which to evaluate risk scores by comparing them against samples of the reference population $\mathcal{P}$.
2.1 Calibration
We say a score function is calibrated over a population $\mathcal{P}$ if and only if for all values $r\in[0,1]$ with $\mathbb{P}\{f_{\theta}(X)=r\}>0$, we have
$\mathbb{P}[Y=1\mid f_{\theta}(x)=r]=r.$ (1)
This condition asks that, over the set of all instances $x$ with score value $r$, an $r$ fraction of those instances must indeed have a positive label. Importantly, calibration is defined with respect to a population, it does not measure a model’s ability to discriminate between instances. A model that outputs the constant value $\mu=\operatorname{\mathbb{E}}[Y]$ on all instances is calibrated, by definition. In particular, we can always achieve calibration by setting $f_{\theta}(x)$ with the average value of $y$ among all instances $x^{\prime}$ in a given partition such that $f_{\theta}(x^{\prime})=f_{\theta}(x).$
We use Expected Calibration Error (ECE) as the primary metric to empirically evaluate calibration. It is defined as the expected absolute difference between a classifier’s confidence in its predictions and the accuracy on the same predictions. More formally, given $n$ triplets $(x_{i},y_{i},r_{i})$ where $r_{i}$ denotes the model’s score value for the corresponding data point $(x_{i},y_{i})$. The ECE is defined as
$\mathrm{ECE}\coloneq\frac{1}{n}\sum_{m}\left|\sum_{i\in B_{m}}y_{i}-\sum_{i\in B_{m}}r_{i}\right|,$ (2)
where data points are grouped into $M$ equally spaced bins $B_{m}$ according to their score values. We use $M=10$ in our evaluation, which is a commonly used value [59]. We also provide a measure of ECE over quantile-based score bins, as well as Brier score [28] to allow for a more complete picture. Furthermore, we use reliability diagrams [30] to aid a visual interpretation of calibration. These diagrams plot expected sample accuracy as a function of confidence. Any deviation from a perfect diagonal represents miscalibration.
We can strengthen the calibration condition in (1) by requiring it in multiple subgroups of the population. Specifically, letting $G$ denote any discrete random variable, we can require the conditional calibration condition $\mathbb{P}[Y=1\mid G=g\,,f_{\theta}(x)=r]=r$ for every setting $g$ of the random variable $G.$ This is often used to define fairness.
2.2 Predictive performance
When solving classification problems it’s common practice to threshold risk scores to obtain a classifier. In our notation this corresponds to thresholding the risk scores $f_{\theta}(x)$:
$c(x)=\mathds{1}\{f_{\theta}(x)>\tau\}.$
The classifier $c$ that minimizes the misclassification error $\operatorname{\mathbb{E}}\mathds{1}\{c(X)\neq Y\}=\mathbb{P}\{c(X)\neq Y\}$ is given by $c^{*}(x)=\mathds{1}\{f^{*}(x)>0.5\}$, where $f^{*}$ is the Bayes optimal scoring function. In the following, when we use accuracy we refer to the fraction of correct predictions after thresholding. If not specified otherwise we use $\tau=0.5$. This corresponds to an argmax operator applied to the class probabilities. It is important to note that a classifier can achieve perfect accuracy even when derived from a suboptimal scoring function. Thus, accuracy alone provides an incomplete picture of a model’s ability to express uncertainty.
Instance ranking.
The area under the receiver operating characteristic curve (AUC) is a rank-based measure of predictive model performance. It measures the probability that a randomly chosen positive observation ($Y=1$) will have a higher score than a randomly chosen negative observation ($Y=0$). A high AUC value in no way reflects accurate or calibrated probability estimates, it relates only to the signal-to-noise ratio in the risk scoring function [11]. Using the AUC metric allows us to neatly separate risk score calibration from their predictive signal, although both are crucial for accurate class predictions.
3 Evaluating language models as risk scores
We are interested in the ability of LLMs to express natural uncertainty in outcomes. Therefore, we construct unrealizable binary prediction tasks, and test the model’s ability to reflect natural variations in underspecified individual outcomes. Specifically, we prompt models with feature values $x$ to elicit risk scores $r$ and then evaluate these scores against ground truth labels $y$ (see Figure 1).
3.1 Prediction tasks
We construct natural language prediction tasks from the American Community Survey (ACS) Public Use Microdata Sample (PUMS). [^3] https://www.census.gov/programs-surveys/acs/microdata.html The data contains survey responses of about 3.2 million anonymous individuals, carefully curated to offer statistical insights into the population of the United States. We refer to the data as Census data. The Census data contains demographic attributes, as well as information related to income, employment, health, transportation, and housing. Prediction tasks are defined by selecting a subset of attributes to define the features and one attribute to be the label. We threshold continuous target variables and bin multi-class predictions to obtain a binary classification task. Specifically, we test models on their ability to reflect natural variations in the outcome across the benchmark population. To enable straightforward comparison with existing tabular benchmarks, we consider natural text analogues to the tasks in the popular folktables benchmark package [9]. Appendix B describes each task in further detail.
Natural uncertainty in risk scoring.
Prediction tasks on human populations typically come with natural outcome uncertainty, meaning that the target label is not uniquely determined by the input features (also known as aleatoric uncertainty [60]). The fewer features are provided the higher the uncertainty in the outcome. We take this to our advantage to systematically evaluate calibration. Namely, to be calibrated, a risk score has to reflect both model uncertainty and uncertainty inherent to the prediction task. In fact, the optimal predictor would often output low-confidence answers. In contrast, prevailing question-answering benchmarks have no data uncertainty, and thus require high confidence for an accurate model to be calibrated. Underspecified, non-realizable prediction tasks allow us to circumvent such potential confounding between calibration and accuracy.
3.2 Extracting risk scores
To extract risk scores from LLMs, we map each inference problem to a natural text prompt. For a given data point, we specify the classification task in the prompt and extract class probabilities from the model’s next token probabilities, similar to traditional question-answering interfaces. The prompt consists of three components (as shown in Figure 1):
  • Instantiating population: We first instantiate the population $\mathcal{P}$ in the prompt context. This corresponds to the population represented by our reference data. We use third-person prompting: “The following data describes a survey respondent. The survey was conducted among US
    residents in 2018. Please answer the question based on the information provided.
    ” This step is typically not needed in supervised classification tasks, because the training data implicitly defines the population. However, for LLMs this is particularly important for evaluating calibration outside the realizable setting, as risk scores cannot in general simultaneously be calibrated to different populations. Skipping this step could provide insights related to alignment [22, 45, 46] rather than calibration.
[…]
When using the constructed multiple-choice prompts, we query the models and extract scores from the next token probabilities for the choice labels A,B as $r_{i}=\mathbb{P}(\texttt{A})/(\mathbb{P}(\texttt{A})+\mathbb{P}(\texttt{B}))$, following the methodology of standard question-answering benchmarks [39, 45]. As LLMs are known to have ordering biases in multiple-choice question-answering [47, 65], we evaluate responses on all choice orderings and average the resulting scores. When using numeric prompting, we prefix the answer with ‘0.’ to improve the likelihood of a direct numeric response, and run two forward passes, selecting the highest likelihood numeric token at each iteration. We refer to Xiong et al. [37] for an overview on alternative design choices on how to elicit confidence scores from LLMs.
3.3 The folktexts package
The folktexts package is designed to offer a flexible interface between tabular prediction tasks and natural language question-answering tasks in order to extract risk scores from LLMs. The package makes available the ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, and ACSTravelTime prediction tasks [9] as natural language benchmarks, together with various functionalities to customize the task definitions (e.g., use a different set of features to predict income) and subsample the reference data (e.g., predict income only among college graduates in California). The set of attributes available to define the features and label can be found in the ACS PUMS data dictionary.444 https://www.census.gov/programs-surveys/acs/microdata/documentation.html. Additionally, folktexts is compatible with open-source models running locally, as well as with closed-source models hosted through a web API. However, APIs must make available the next token probabilities for each forward pass instead of returning discrete text completions — the OpenAI API is one example that is compatible with folktexts out of the box.
In addition to providing a reproducible way to extract risk scores from LLMs, folktexts also offers pre-implemented evaluation metrics to benchmark and compare the calibration and accuracy of LLMs, as well as easy plotting of group-conditional calibration curves for a cursory view of potential biases. The package is easy to use within a python notebook, as a dependency, or directly from the command line. Further details on usage and design choices are available in Appendix C.
4 Empirical findings
We use folktexts to evaluate several recently released models together with their instruction-tuned counterparts: the Llama 3 models [66], including the 8B and the 70B versions, the Mistral 7B [67], the Mixtral 8x7B and 8x22B variants [68], the Yi 34B [69], and the Gemma [70] 2B and 7B variants. We also evaluate GPT 4o mini [71] through the OpenAI API (note that no base model version is available). Instruction-tuned models are marked ‘(i.t.)’. For comparison, results are also shown for a logistic regression (LR) model, and a gradient boosted decision trees model (XGBoost). The XGBoost model [72] is generally regarded as the state-of-the-art in tabular data tasks [73].
In this section we focus on the ACSIncome prediction task, which is the default folktexts benchmarking task. It consists in predicting whether a person’s income is above or below \$50K from 10 demographic features, and closely emulates the popular UCI Adult prediction task [74]. The evaluation test set consists of 160K randomly selected samples from the 2018 Census data. A separate set of 1.5M samples is used to train the supervised LR and XGBoost models, LLMs are used as zero-shot classifiers without fine-tuning. Both multiple-choice prompting and chat-style numeric prompting were used to obtain two separate risk score distributions for each model (as described in Section 3.2). We focus on analyzing multiple-choice prompting results, as it is arguably the standard in LLM benchmarking [19, 20, 21, 22, 23, 24, 39, 45, 46, 47, 48]. Appendix A presents additional results and plots for the experiments analyzed in this section, including a more in-depth look into numeric prompting results, as well as results on alternative prediction tasks. Experiments were ran on a cluster with Nvidia-A100-80GB GPUs, consuming an approximate total of 500 GPU hours.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.24 0.24 0.85 0.74 0.05 0.16 0.83 0.78
Mixtral 8x22B (it) 0.21 0.22 0.85 0.76 0.11 0.17 0.84 0.77
Mixtral 8x22B 0.17 0.19 0.85 0.68 0.13 0.18 0.82 0.74
Llama 3 70B (it) 0.27 0.27 0.86 0.69 0.25 0.23 0.84 0.67
Llama 3 70B 0.20 0.20 0.86 0.70 0.27 0.24 0.82 0.54
Mixtral 8x7B (it) 0.16 0.18 0.86 0.78 0.10 0.17 0.84 0.76
Mixtral 8x7B 0.17 0.21 0.83 0.65 0.07 0.17 0.81 0.78
Yi 34B (it) 0.19 0.19 0.86 0.72 0.22 0.21 0.80 0.48
Yi 34B 0.25 0.22 0.85 0.62 0.15 0.19 0.83 0.61
Llama 3 8B (it) 0.32 0.30 0.85 0.62 0.23 0.23 0.81 0.67
Llama 3 8B 0.25 0.26 0.81 0.38 0.14 0.24 0.63 0.40
Mistral 7B (it) 0.21 0.22 0.83 0.77 0.16 0.19 0.83 0.70
Mistral 7B 0.20 0.23 0.80 0.73 0.36 0.32 0.75 0.49
Gemma 7B (it) 0.61 0.59 0.84 0.37 0.33 0.30 0.78 0.42
Gemma 7B 0.24 0.27 0.76 0.37 0.15 0.20 0.80 0.73
Gemma 2B (it) 0.63 0.63 0.73 0.37 0.28 0.31 0.50 0.37
Gemma 2B 0.14 0.25 0.62 0.45 0.37 0.37 0.50 0.63
LR 0.03 0.18 0.79 0.74 0.03 0.18 0.79 0.74
XGBoost 0.00 0.13 0.90 0.82 0.00 0.13 0.90 0.82
Table 1: Zero-shot LLM results on the ACSIncome benchmark task, together with results for LR and XGBoost baselines fitted on 1.5M samples. Table cells are colored using a continuous color map between the worst ( in orange) and best ( in cyan) results of each column. Models generally achieve high predictive signal (high AUC) but poor calibration (high ECE). Numeric prompting leads to improved ECE but worse AUC.
4.1 Benchmark results
We perform a comprehensive evaluation of risk scores output by LLMs along multiple metrics. Results are summarized in Table 1.
Multiple-choice prompting.
We observe that a majority of LLMs (all of size 8B or larger) outperform the linear model baseline (LR) in terms of predictive power (AUC). However, LLMs are clearly far from matching the supervised baselines with respect to calibration: all language models achieve very high calibration error (ECE), while baselines achieve near-perfect calibration. Due to this high miscalibration, models struggle to translate scores with high predictive signal into high accuracy. In fact, while most models achieve high AUC, most models struggle to surpass the supervised linear baseline in terms of accuracy. We recall that AUC is agnostic to calibration, while accuracy on the maximum likelihood answer is not. All of the Gemma models have worse than random accuracy, despite having clearly above random AUC (random would be 0.5). Only the instruction-tuned Mistral models (7B, 8x7B, and 8x22B) outperform the linear model in terms of accuracy. Interestingly, while larger models achieve higher AUC, calibration is not reliably improved by model size — differences across model families are more pronounced than across model sizes.
Finally, we focus on comparing base models to their instruction-tuned counterparts, marked with ‘(it)’. A striking trend is visible across the board: instruction-tuning generally worsens calibration (higher ECE) when using multiple-choice prompting. At the same time, we generally see improvements in AUC and accuracy after instruction-tuning. Appendix A.2 presents results on the four additional prediction tasks. The same trend of instruction-tuning leading to worse calibration and higher AUC is broadly replicated. However, performance across different tasks is somewhat inconsistent: LLMs generally have stronger predictive signal than a supervised linear model on the income prediction and travel-time prediction tasks, but consistently underperform the linear baseline on the address change task (ACSMobility).
Numeric prompting.
We observe broad improvements in calibration (lower ECE) and Brier score loss across instruction-tuned models when using numeric prompting. Figure 2 shows the change in ECE between both prompting schemes. The trend is clear across most instruct models across all five benchmark tasks. The Yi 34B model is the exception, showing outlier results throughout the different experiments we conduct. On the other hand, results for base models are less conclusive (see Figure A1). At the same time, numeric prompting leads to small but consistent drops in predictive power of risk scores (AUC) on 4 out of 5 benchmark tasks, including ACSIncome. Figure A2 shows these changes visually for all tasks. Finally, we point out that the GPT 4o mini model produces a surprisingly well-calibrated risk score distribution for the income prediction task ($\text{ECE}=0.05$). In fact, for each benchmark task, at least one LLM is able to produce a remarkably well-calibrated score distribution (although a different model for different tasks). This promising result points to the fact that some small amount of data will likely always be needed to properly evaluate LLMs capabilities to model human population statistics, but such modelling can often be done with high degree of confidence and a good understanding on which outputs might be wrong.
Figure 2: Change in calibration error (ECE) of instruction-tuned models when using numeric risk prompting (orange circles) versus multiple-choice prompting (blue squares). Improvement/deterioration is represented by green/red arrows, respectively. An overwhelming majority of model/task pairs see calibration improvements.
Figure 3: Calibration curves for base and instruction-tuned versions of the largest models studied, on the ACSIncome task. Curves are computed using $10$ quantile-based score bins. Risk scores were generated using multiple-choice-style prompting (left plots) or numeric chat-style prompting (right plots).
4.2 Score distribution
To get additional insights into the difference between base models and their instruction-tuned counterparts, we inspect their risk score distributions more closely.
Figure 3 shows calibration curves of the largest LLMs studied, for both base and instruct variants and for both prompting schemes. When using multiple-choice prompting (left-most plots of Fig. 3), both base and instruction-tuned models have poor score calibration, but failure modes are entirely different: Base models output under-confident scores, while instruction-tuned models output over-confident scores. To quantify this result, we introduce a measure of risk score confidence bias similar to the ECE metric:
$\mathrm{R}_{\mathrm{bias}}\coloneq\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left[\mathrm{Conf}(B_{m})-\mathrm{Acc}(B_{m})\right],$
where $M$ is the number of score bins, $B_{m}$ is the set of samples in score bin $m$, $\mathrm{Acc}(.)$ is the accuracy on a given set of samples, and $\mathrm{Conf}(.)$ is the confidence on a given set of samples measured as the mean risk score for the highest likelihood class. Figure 4(a) shows the risk score confidence bias results. The two miscalibration modes are evident: confidence bias is higher for instruction-tuned models and lower (or even negative) for base models. That is, instruction-tuned models are generally over-confident in their predictions, outputting higher scores than their accuracy would warrant, while no such trend is visible for base models. On the other hand, when using numeric prompting, the two aforementioned failure modes are no longer evident, and the differences between base and instruction-tuned models are blurred (right-most plots of Fig. 3). In fact, Figure 4(b) shows a trend reversal when using numeric prompting: Base models now show a higher over-confidence in their risk scores, while instruction-tuned models show approximately neutral score bias.
(a) Multiple-choice prompting.
(b) Numeric prompting.
Figure 4: Risk score confidence bias for all LLMs on the ACSIncome task. Negative values indicate under-confident risk scores (overestimating uncertainty), while positive values indicate over-confident risk scores (underestimating uncertainty). Instruction-tuned models are generally over-confident when using multiple-choice prompting (Fig. 4(a)), but this bias is considerably diminished when using numeric prompting (Fig. 4(b)).
Figure 5 shows the risk score distribution for a variety of model pairs, produced using multiple-choice prompting. The score distributions for base and instruct model variants are immediately distinguishable: base models consistently produce low-variance distributions centered around 0.5, while instruction-tuned variants often output scores near 0 or 1. The same trend is visible on all model pairs, with Yi 34B showing the smallest difference between base and instruct variants. The score distributions produced by base/i.t. model pairs are markedly different, even among base/i.t. pairs achieving the exact same AUC; e.g., Llama 3 70B and Mixtral 8x22B. For the largest Llama (70B) and largest Mistral (8x22B) models, no predictive performance is gained by instruction tuning, but answers of the instruction-tuned models have significantly higher (over-)confidence and worse calibration. In fact, while the base Llama 3 70B has an average of 0.07 under-confidence bias, the instruct variant produces risk scores on average 0.22 over-confident (see Figure 4(a)). Note that the instruction-tuned Gemma models degenerate into predicting only positive outcomes with high confidence score, hence why they have the lowest accuracy, worst Brier score, and worst ECE.
Crucially, our evaluation reveals a previously unreported shortcoming of using multiple-choice prompting with instruction-tuned models: instruction-tuning polarizes score distributions, even if the true outcome has high entropy. Standard realizable knowledge testing benchmarks can easily disguise this polarization phenomenon as improper quantification of model uncertainty. In fact, it seems evident that it is improper quantification of uncertainty in general, regardless of underlying uncertainty in the modelled distribution. The following subsection goes further in-depth on the influence of data uncertainty in score distribution. Appendix A.1 analyzes numeric prompting results.
Figure 5: Risk score distribution for base and instruction-tuned model pairs on the ACSIncome task, using multiple-choice prompting. After instruction-tuning, models exhibit high confidence, but worse calibration in general. The XGBoost scores showcase a perfectly calibrated distribution ($\text{ECE}\approx 0.00$).
4.3 Varying degree of uncertainty
Next, we consider the dependence of multiple-choice risk scores on the available evidence. For this study we use the Mixtral 8x7B model, which achieves the best (lowest) Brier score among evaluated models (reflecting both high accuracy and high calibration). We compute income prediction risk scores with increasing evidence: starting with only 2 features, and iteratively adding 2 features at a time (see results in Figure A7). This sequence demonstrates how unrealizable, underspecified prediction tasks differ from realizable prediction tasks. Predicting income based on an individual’s place of birth (POBP) and race (RAC1P) is naturally not possible to a high degree of accuracy, forcing any calibrated model to output lower-confidence risk scores. Indeed, both base and instruct variants correctly output lower confidence scores for the smaller feature sets when compared with the larger feature sets (compare left-most to right-most plots of Fig. A7). However, instruction-tuning still leads to a clear polarization of risk score distribution, regardless of true data uncertainty: Score variance for base models is in range $\sigma\in[0.02,0.06]$, while for i.t. models it’s in range $\sigma\in[0.16,0.41]$. Appendix A.4 goes further in-depth on how score distributions change with varying data uncertainty.
In addition to providing insights into risk scores, varying individual features in the prediction task also provides insights into LLM feature importance. Appendix A.5 presents LLM feature importance results and discusses the main differences to traditional supervised learning models.
Figure 6: Calibration curves for the Mixtral 8x22B (top) and the Yi 34B (bottom) models, across different race subpopulations on the ACSIncome task. Left: Multiple-choice-style prompting. Right: Numeric or chat-style prompting. The models shown are the worst offenders in terms of subgroup-specific score bias. On average, Black individuals see a lower score ($x$ axis) for the same true probability of positive outcomes ($y$ axis).
4.4 Subgroup calibration
Finally, we evaluate risk score calibration across different subpopulations, such as typically done as part of a fairness audit. Figure 6 shows calibration curves for two sets of models on the ACSIncome task, evaluated on three subgroups specified by the instantiation of the race attribute in the data. We pick the three race categories with the largest representation in the ACS data. Note that a positive prediction of $Y=1$ is arguably the advantageous outcome, as it corresponds to the high-income category (“Earns above \$50,000 per year”). On the two models shown, samples belonging to the ‘Black’ population see consistently lower scores for the same positive label probability when compared to the ‘Asian’ or ‘White’ populations. The Mixtral 8x22B (it) and Yi 34B (it) models shown are the worst offenders, with an average risk score difference between Asian and Black groups of 0.17 and 0.13 respectively; i.e., a Black individual will receive on average a 0.17 or 0.13 lower score than an Asian individual for the same true probability of high-income $\mathbb{P}\{Y=1\}$. This poses a higher bar for Black individuals to get a positive prediction of “high-income”. Note that the remaining models studied do not show such striking differences in group-conditional calibration. In fact, this score bias can be reversed for some base models, overestimating scores from Black individuals compared with other subgroups (results for all models shown in Appendix A.3). Such differences in score calibrations arguably warrant a more in-depth analysis that escapes the scope of this paper. We raise concerns regarding subgroup miscalibration, which should caution practitioners against using such scores in consequential domains without a comprehensive fairness audit.
5 Discussion
We introduced folktexts, a software package that provides datasets and tools to evaluate risk scores produced by language models. Unlike most existing LLM benchmarks, the datasets we introduced have inherent outcome uncertainty, making them useful as a basis for systematic evaluation of uncertainty quantification in LLMs. While uncertainty on realizable tasks reflects only model uncertainty (i.e., whether the model is aware of its lack of knowledge), uncertainty on unrealizable tasks is itself a type of knowledge over the underlying data distribution.
Our empirical findings show that LLM risk scores produced using standard multiple-choice Q&A generally have strong predictive signal, but are wildly miscalibrated. Such models may be good for knowledge testing, but lack adequate indicators of uncertainty, making them unsuitable for synthetic data generation. Instruction-tuned models generally have worse calibration but slightly higher accuracy and AUC than their base-model counterparts on multiple-choice answers. In fact, instruction-tuning leads to marked polarization of multiple-choice answer distribution, regardless of underlying true data uncertainty. This reveals a general inability of instruction-tuned LLMs to quantify uncertainty using multiple-choice Q&A. At the same time, verbally querying models for numeric probability estimates considerably improves calibration of instruction-tuned models, but at a small cost in AUC. Going forward, we envision that future package extensions will include other uncertainty quantification methods, such as confidence intervals, or conformal prediction methods.
Acknowledgments
We thank Florian Dorner, Mila Gorecki, and Ricardo Dominguez-Olmedo for invaluable feedback on an earlier version of this paper. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting André F. Cruz.
Supplementary Materials A Additional results
This appendix section shows additional results and corresponding plots to support the insights presented in Section 4. Section A.1 shows results using a chat-style verbalized numeric prompting scheme. Section A.2 shows results on four extra benchmark tasks made available with folktexts. Section A.3 extends the discussion on subgroup calibration and algorithmic fairness on the ACSIncome task. Section A.4 goes further in-depth on how to use folktexts to control data uncertainty in the benchmark prediction tasks. Finally, Section A.5 presents and discusses results on feature importance for LLM predictions.
A.1 Additional results using verbalized numeric prompting
Figure A1 shows the change in calibration error (ECE) between using multiple-choice prompting and verbalized numeric prompting, on all five benchmark tasks. Instruction-tuned models (top rows) show ECE improvements on an overwhelming majority of model/task pairs, while base models (bottom rows) show less consistent results. However, using numeric prompting comes at a consistent cost of diminished predictive power (AUC) of the risk scores, shown in Figure A2. A majority of model/task pairs have worse AUC with numeric prompting, with the exception of the employment prediction task. One potential explanation for this generalized decrease in AUC lies in the fact that numeric prompting generates a large number of tied risk scores. Figure A3 shows the score distribution of all models using numeric prompting (compare with multiple-choice prompting shown in Figure 5). While multiple-choice prompting produces a smoother continuous score distribution, numeric prompting generally results in a small set of possible uncertainty estimates. This arguably makes intuitive sense, as numeric prompting produces uncertainty estimates in discrete token space, while multiple-choice prompting produces uncertainty estimates in the continuous token-probability space.
A.2 Results on additional benchmark tasks
The main body of the paper focuses on results on the ACSIncome prediction task. This task is arguably the most popular for benchmarking on tabular data, as it closely mirrors the older but widely used UCI Adult dataset [9, 74]. The folktexts package makes available natural-language versions of four additional tabular data tasks: ACSEmployment, ACSMobility, ACSTravelTime, and ACSPublicCoverage.
Tables A1A4 show results for the ACSEmployment, ACSMobility, ACSTravelTime, and ACSPublicCoverage tasks, respectively. Trends discussed in the main body of the paper are broadly confirmed. Models’ moderate predictive performance is accompanied by substantial miscalibration of their risk scores. Additionally, base models output low-variance high-uncertainty score distributions, while instruction-tuned models output high-variance low-uncertainty score distributions. There is clear predictive signal for large models across all tasks (e.g., see the AUC of Llama 3 70B it, or Mixtral 8x22B it). However, the extent to which models’ scores are predictable varies substantially across tasks. Most models surpass the AUC of the linear baseline (LR) on the ACSTravelTime task, as well as the main ACSIncome task; but consistently lag behind the linear baseline on the ACSMobility task. On the ACSEmployment and ACSPublicCoverage tasks, the best performing models manage to match the linear baseline AUC.
These findings pose into question one of the main advantages of using LLMs for risk scoring: the fact that no labeled data is required. Given the inconsistency of model performance, some small amount of testing data may always be needed to assert reliability of results.
Figure A1: Change in calibration error (ECE) when using numeric risk prompting ($\bullet$) versus multiple-choice prompting ($\blacksquare$). Instruction-tuned models (top rows) show substantial calibration improvements, while base models (bottom rows) show mixed results. Green/red arrows signal ECE improvement/degradation.
Figure A2: Change in predictive power (AUC) when using numeric risk prompting ($\bullet$) versus multiple-choice prompting ($\blacksquare$). Both instruction-tuned models (top rows) and base models (bottom rows) generally achieve worse AUC with numeric prompting. Green/red arrows signal AUC improvement/degradation, respectively.
Figure A3: Distribution of risk scores produced using numeric prompting on the ACSIncome benchmark task. A baseline score distribution that achieves $0.00$ calibration error is shown in green (XGBoost model). For each model, risk scores produced in this manner fall into only a few different possible values, contrasting with the neatly continuous distribution produced by multiple-choice prompting. This fact leads to numerous more ties among predicted risk scores, which can explain the reduced AUC performance with this prompting scheme. Nonetheless, calibration error is considerably smaller for instruction-tuned models.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.28 0.29 0.79 0.65 0.23 0.23 0.80 0.73
Mixtral 8x22B (it) 0.38 0.39 0.60 0.51 0.06 0.14 0.87 0.79
Mixtral 8x22B 0.21 0.24 0.86 0.52 0.15 0.18 0.82 0.80
Llama 3 70B (it) 0.17 0.19 0.85 0.73 0.05 0.14 0.88 0.81
Llama 3 70B 0.25 0.26 0.82 0.52 0.05 0.15 0.86 0.78
Mixtral 8x7B (it) 0.22 0.24 0.82 0.73 0.07 0.15 0.87 0.78
Mixtral 8x7B 0.30 0.31 0.81 0.45 0.08 0.17 0.81 0.73
Yi 34B (it) 0.14 0.21 0.79 0.69 0.15 0.21 0.81 0.51
Yi 34B 0.08 0.23 0.70 0.62 0.13 0.23 0.66 0.50
Llama 3 8B (it) 0.07 0.19 0.79 0.74 0.08 0.17 0.82 0.77
Llama 3 8B 0.34 0.34 0.76 0.45 0.15 0.23 0.75 0.46
Mistral 7B (it) 0.35 0.36 0.72 0.63 0.04 0.19 0.79 0.69
Mistral 7B 0.26 0.30 0.76 0.45 0.14 0.19 0.80 0.79
Gemma 7B (it) 0.36 0.38 0.59 0.58 0.04 0.22 0.71 0.60
Gemma 7B 0.15 0.25 0.65 0.48 0.35 0.38 0.50 0.51
Gemma 2B (it) 0.38 0.41 0.42 0.46 0.12 0.27 0.46 0.46
Gemma 2B 0.01 0.24 0.63 0.54 0.01 0.23 0.57 0.53
LR 0.02 0.15 0.86 0.78 0.02 0.15 0.86 0.78
XGBoost 0.00 0.12 0.91 0.83 0.00 0.12 0.91 0.83
Table A1: Zero-shot LLM results on the ACSEmployment benchmark task, together with supervised learning baselines fitted on 2.9M samples.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.26 0.26 0.57 0.73 0.22 0.25 0.49 0.73
Mixtral 8x22B (it) 0.40 0.40 0.51 0.39 0.05 0.20 0.54 0.73
Mixtral 8x22B 0.11 0.21 0.55 0.73 0.13 0.22 0.49 0.73
Llama 3 70B (it) 0.20 0.25 0.57 0.58 0.05 0.20 0.52 0.73
Llama 3 70B 0.22 0.24 0.55 0.53 0.06 0.20 0.53 0.73
Mixtral 8x7B (it) 0.26 0.26 0.58 0.73 0.11 0.21 0.51 0.73
Mixtral 8x7B 0.14 0.21 0.57 0.73 0.24 0.25 0.48 0.73
Yi 34B (it) 0.09 0.20 0.56 0.72 0.23 0.25 0.50 0.27
Yi 34B 0.07 0.20 0.57 0.73 0.15 0.23 0.52 0.44
Llama 3 8B (it) 0.15 0.22 0.56 0.70 0.11 0.21 0.49 0.73
Llama 3 8B 0.10 0.20 0.55 0.73 0.14 0.21 0.51 0.72
Mistral 7B (it) 0.26 0.26 0.57 0.73 0.17 0.23 0.49 0.73
Mistral 7B 0.20 0.23 0.53 0.73 0.27 0.27 0.50 0.73
Gemma 7B (it) 0.25 0.26 0.58 0.73 0.25 0.26 0.49 0.73
Gemma 7B 0.41 0.37 0.50 0.27 0.19 0.24 0.49 0.73
Gemma 2B (it) 0.73 0.73 0.52 0.27 0.02 0.20 0.50 0.73
Gemma 2B 0.25 0.26 0.51 0.34 0.27 0.27 0.50 0.73
LR 0.02 0.19 0.61 0.74 0.02 0.19 0.61 0.74
XGBoost 0.00 0.16 0.74 0.76 0.00 0.16 0.74 0.76
Table A2: Zero-shot LLM results on the ACSMobility benchmark task, together with supervised learning baselines fitted on 0.6M samples.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.39 0.40 0.65 0.55 0.15 0.27 0.58 0.57
Mixtral 8x22B (it) 0.31 0.33 0.66 0.59 0.12 0.24 0.64 0.59
Mixtral 8x22B 0.20 0.28 0.63 0.44 0.30 0.34 0.57 0.58
Llama 3 70B (it) 0.15 0.24 0.70 0.60 0.12 0.24 0.64 0.53
Llama 3 70B 0.09 0.24 0.67 0.55 0.08 0.25 0.52 0.46
Mixtral 8x7B (it) 0.45 0.45 0.66 0.52 0.09 0.24 0.61 0.57
Mixtral 8x7B 0.28 0.32 0.60 0.44 0.07 0.25 0.57 0.58
Yi 34B (it) 0.35 0.36 0.65 0.56 0.06 0.25 0.50 0.44
Yi 34B 0.08 0.24 0.62 0.56 0.14 0.27 0.53 0.44
Llama 3 8B (it) 0.19 0.28 0.60 0.57 0.11 0.25 0.56 0.56
Llama 3 8B 0.08 0.25 0.53 0.56 0.12 0.26 0.48 0.44
Mistral 7B (it) 0.41 0.42 0.59 0.57 0.11 0.25 0.55 0.56
Mistral 7B 0.05 0.25 0.57 0.56 0.44 0.44 0.50 0.56
Gemma 7B (it) 0.42 0.43 0.53 0.56 0.10 0.26 0.49 0.44
Gemma 7B 0.04 0.24 0.61 0.58 0.03 0.25 0.52 0.55
Gemma 2B (it) 0.34 0.36 0.49 0.56 0.19 0.28 0.50 0.56
Gemma 2B 0.09 0.26 0.48 0.44 0.44 0.44 0.50 0.56
LR 0.04 0.24 0.58 0.56 0.04 0.24 0.58 0.56
XGBoost 0.02 0.19 0.77 0.70 0.02 0.19 0.77 0.70
Table A3: Zero-shot LLM results on the ACSTravelTime benchmark task, together with supervised learning baselines fitted on 1.3M samples.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.33 0.34 0.71 0.60 0.10 0.20 0.68 0.73
Mixtral 8x22B (it) 0.24 0.25 0.70 0.72 0.04 0.18 0.71 0.75
Mixtral 8x22B 0.32 0.30 0.59 0.30 0.29 0.29 0.54 0.70
Llama 3 70B (it) 0.16 0.21 0.69 0.75 0.13 0.20 0.73 0.75
Llama 3 70B 0.18 0.22 0.67 0.63 0.12 0.21 0.64 0.53
Mixtral 8x7B (it) 0.20 0.23 0.70 0.74 0.06 0.19 0.69 0.74
Mixtral 8x7B 0.41 0.37 0.57 0.30 0.20 0.25 0.56 0.70
Yi 34B (it) 0.06 0.19 0.67 0.74 0.22 0.24 0.57 0.31
Yi 34B 0.04 0.21 0.59 0.70 0.09 0.20 0.67 0.64
Llama 3 8B (it) 0.11 0.21 0.59 0.71 0.17 0.22 0.64 0.68
Llama 3 8B 0.41 0.38 0.55 0.30 0.20 0.25 0.51 0.34
Mistral 7B (it) 0.30 0.30 0.61 0.70 0.07 0.20 0.67 0.65
Mistral 7B 0.29 0.30 0.45 0.30 0.30 0.30 0.50 0.70
Gemma 7B (it) 0.30 0.34 0.46 0.50 0.18 0.24 0.57 0.61
Gemma 7B 0.15 0.23 0.49 0.49 0.18 0.26 0.48 0.70
Gemma 2B (it) 0.70 0.70 0.54 0.30 0.24 0.29 0.42 0.42
Gemma 2B 0.26 0.28 0.54 0.30 0.30 0.30 0.50 0.70
LR 0.03 0.19 0.70 0.72 0.03 0.19 0.70 0.72
XGBoost 0.00 0.14 0.84 0.80 0.00 0.14 0.84 0.80
Table A4: Zero-shot LLM results on the ACSPublicCoverage benchmark task, together with supervised learning baselines fitted on 1.0M samples.
(a) Multiple-choice prompting.
(b) Numeric prompting.
Figure A4: Calibration curves for the ‘GPT 4o mini’ model, across different race sub-populations on the ACSIncome task, computed using $10$ quantile-based bins. This model is only available through a web API, and no base model variant is available. Numeric prompting (Fig. 4(b)) leads to reduced differences in group-conditional calibration (improved group fairness), as well as large improvements in overall calibration.
Figure A5: Calibration curves across different race sub-populations on the ACSIncome task, computed using $10$ quantile-based score bins, with $95\%$ confidence intervals. Using multiple-choice prompting to generate risk scores (left plots) leads to sizeable calibration error for both base and instruction-tuned models. The models shown in this figure generally do not follow the same group bias trend seen in the Mixtral 8x22B and Yi 34B models (Figure 6).
A.3 Additional subgroup calibration results
This section contains additional subgroup calibration plots on the ACSIncome task. Figure 6 shows the 2 worst offenders: Mixtral 8x22B and Yi 34B. Figures A4 and A5 show the remaining models. The same trend is visible but to a lesser extent: Instruction-tuned models show a bias towards lower scores for Black individuals, even after controlling for label prevalence.
We can quantify the positive score bias by evaluating the signed calibration error (SCE):
$\mathrm{SCE}\coloneq\frac{1}{n}\sum_{m=1}^{M}\sum_{i\in B_{m}}(r_{i}-y_{i}),$ (3)
where $M$ is the number of score buckets, $B_{m}$ is the set of sample indices belonging to bucket $m$, $r_{i}=f_{\theta}(x_{i})$ is the risk score given to sample $x_{i}$ and $y_{i}$ is its label. This metric does not evaluate overall calibration, as a value of 0 does not indicate a calibrated classifier. Instead, negative/positive values indicate a bias towards lower/higher risk scores, respectively. If higher scores are related to positive real-world outcomes (e.g., when predicting income for a loan application), then a bias towards lower scores on samples of specific protected subgroups would likely lead to unfair real-world outcomes. Conversely, if a negative class prediction is associated with a positive outcome (e.g., when predicting risk of recidivism), then a bias towards lower scores would be beneficial for the affected group.
Figure A6 shows the difference between the signed calibration error (SCE) on different group pairings, on the ACSIncome task. Positive predictions ($Y=1$) correspond to the advantaged high-income class. Negative differences, $\Delta_{SCE}<0$, indicates advantaged scores for Black individuals, while positive differences, $\Delta_{SCE}>0$, indicate disadvantaged scores for Black individuals. Interestingly, while subgroup calibration curves appear similar for base models and disadvantage Black individuals for instruction-tuned models (see Figures 6A4, and A5), this is not entirely reflected on the SCE metric. Indeed, a clear-cut split is visible: base models benefit the score of Black individuals, and instruct models disadvantage the score of Black individuals. This finding could be partially explained by the fact that base models produce score distributions with low variance, and instruct models produce high-variance polarized outcomes. Specifically, two conclusions can be drawn from the score distributions produced by base models: (1) models under-estimate the score of high-income earners (which are disproportionately Asian), and (2) models over-estimate the score of low-income earners (which are disproportionately Black). This could lead base models to over-estimate the earnings of the Black population disproportionately to other groups. The opposite is true for instruction-tuned models: high-income earners see their score over-estimated, which benefits groups with a higher prevalence of high earners.
This in no way serves as an exhaustive analysis of risk score fairness on LLMs, as it is bound to be highly task dependent and language model dependent. We simply surface the fact that, on this high-income prediction task, risk scores are not group-calibrated [8] and could lead to unfair outcomes. Crucially, even though race has the lowest mean feature importance among all features (see Appendix A.5), we report and explain how different trends in risk score distributions can effectively lead to unfair outcomes.
Figure A6: Racial group bias in risk score calibration error. When comparing score bias of group $A$ with score bias of group $B$, $\Delta_{\mathrm{SCE}}=\mathrm{SCE}_{A}-\mathrm{SCE}_{B}$, positive values indicate an undue score advantage of group $A$, and negative values an undue score advantage of group $B$. Left: Difference between White and Black groups, $\mathrm{SCE}_{White}-\mathrm{SCE}_{Black}$. Right: Difference between Asian and Black groups, $\mathrm{SCE}_{Asian}-\mathrm{SCE}_{Black}$. Note that the Gemma models were omitted, as their instruct versions degenerate into strictly predicting the same outcome for all samples. Consequently, the two instruct Gemma models are the only exceptions to the trend shown in these plots.
A.4 Varying uncertainty
Figure A7: Shift of score distribution with increasing evidence for the Mixtral 8x7B model (which achieved the best Brier score), using multiple-choice Q&A on the ACSIncome task. Features are described in Table A5. Score distribution gets more discriminative as more evidence is added, successfully increasing scores’ predictive signal (AUC). The true label prevalence is $\mathbb{P}[Y=1]=0.37$.
Figure A8: Evaluation of calibration (ECE) and predictive performance (AUC) on Llama and Mistral models, with an increasing number of features provided on the ACSIncome task. For each dot along the line we add two features, up to all 10 features being used in the point marked with a star. Top row: base models. Bottom row: instruction-tuned models. Models can successfully use each extra feature to increase predictive signal. Calibration trends worse the more features are added for base models, while instruct models show no clear trend.
A simple API call in our package allows for selecting different subsets of attributes to include as features when using the LLM as a predictor. Figure A8 inspects the effect of increasing the feature on models’ calibration and predictive power. Each dot along the line represents an increasing feature set used for LLM predictions, added in order of mean feature importance on all models. Appendix B details all features used in the ACSIncome task. We refer the reader to Ding et al. [9] and the ACS codebook [^5] See the ACS PUMS data dictionary for the full list of available variables:
https://www.census.gov/programs-surveys/acs/microdata/documentation.html
for an in-depth description of each categorical value a feature can take. Predictive signal (ROC AUC) reliably increases with each added feature, for all tested models except the Gemma 2B variants. This is expected with standard supervised learning algorithms trained and evaluated on i.i.d. data, but arguably somewhat unexpected of pre-trained models trained on a variety of datasets that are out-of-distribution relative to the evaluation set. On the other hand, there is no clear trend for calibration: for Mistral models, it seems that calibration actually worsens for larger feature sets, while for Llama models calibration is approximately stable across all points.
This experiment show-cases one unique way of using LLMs with survey prediction tasks: while supervised learning models would have to be retrained every time a different feature set is used, LLMs can freely change the evidence they use to make a prediction. If a model were to exhibit properties of a joint distribution with the ability to marginalize over hidden features, then calibration with respect to evidence $\mathcal{X}$ implies that it is also calibrated with respect to restricted evidence $\mathcal{X}^{\prime}\subset\mathcal{X}$. To what extent a model satisfies such properties is an interesting question for future work; we hope our package proves useful as an investigative tool.
A.5 Feature importance
In this section, we present feature importance results for different LLMs on the ACSIncome prediction task. The importance value of feature $j$ is computed as the drop in AUC after permuting all values of feature $j$ across the dataset. That is, each sample $x$, sees its value for feature $j$ randomly permuted with another sample. This is a common feature importance implementation [75], as it does not rely on any internal characteristics of the model.
Figure A9 shows feature importance values for the largest language models studied (above 40B active parameters). Results for the XGBoost model are also shown in green. Note that XGBoost achieves the best result on every single metric in Table 1. While for supervised models, a given categorical value is nothing more than a 1 or 0, LLMs have the potential to surface the real-world meaning of such values, benefiting from the rich embedding representations of each category. As such, we’d expect to see LLMs assigning higher importance to categorical features. Indeed, Llama 3 models assign considerably higher importance to the occupation feature (OCCP), which is a numerically encoded categorical feature with over 500 different possible values. Conversely, the XGBoost model assigns considerably higher importance than LLMs to ‘work-hours per week’ (WKHP) and ‘age’ (AGEP), both integer-encoded features. Lastly, feature importance results indicate that the studied LLMs do not explicitly use sensitive categories such as age (AGEP), sex (SEX), or race (RAC1P) for risk score estimation.
Interestingly, feature importance is similar for base and instruct variants of the same model. This contrasts with the score distribution and calibration curve results, where all base models followed a similar trend, distinct from their instruction-tuned versions.
Figure A9: Feature importance among the largest language models tested, plus results for the XGBoost baseline. Feature importance values are calculated as the loss in AUC when the values of a given column are randomly permuted [75].
Supplementary Materials B Details on provided benchmark tasks
The folktexts package defines natural-text mappings for a variety of columns in the ACS PUMS data files. Table A5 lists and describes each implemented column-to-text mapping. Any combination of column-to-text objects can be used to create a prediction task from ACS data, both as features and as the prediction target. To enable straightforward comparison with existing benchmarks, we mimic the feature set and population filters used by the prediction tasks available in the popular folktables benchmark package [9]. Specifically, we put forth natural-text variants of the ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, and ACSTravelTime tasks. These prediction tasks define a restricted set of columns from the ACS PUMS data files to be used as input features for machine learning models, as well as a binarized target column. As such, we extend the use of these ACS prediction tasks to benchmark language models, enabling direct comparison with a wide-ranging set of literature works. Although any ACS survey year could be used for benchmarking, we define the standard set of benchmark tasks as those using data from the 2018 1-year-horizon person-level survey (following Ding et al. [9]). Notwithstanding, we welcome the addition of new column-to-text mappings by new users of the package, both for ACS data and for new datasets. The following paragraphs detail each pre-implemented prediction task.
ACSIncome
The goal of the ACSIncome task is to predict whether a person’s yearly income is above \$50,000, given by the PINCP column. The ACS columns used as features are: AGEP, COW, SCHL, MAR, OCCP, POBP, RELP, WKHP, SEX, and RAC1P. The sub-population over which the task is conducted is employed US residents with age greater than 16 years. The ACSIncome prediction task was put-forth as the successor to the popular UCI Adult dataset [74], used extensively in the algorithmic fairness literature. This task is the default task when running the folktexts benchmark.
ACSPublicCoverage
The goal of the ACSPublicCoverage task is to predict whether an individual is covered by public health insurance, given by the PUBCOV column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, CIT, MIG, MIL, ANC, NATIVITY, DEAR, DEYE, DREM, PINCP, ESR, ST, FER, and RAC1P. The sub-population over which the task is conducted is US residents with age below 65 years old, and with personal income below \$30,000.
ACSMobility
The goal of the ACSMobility task is to predict whether an individual has changed their home address in the last year, given by the MIG column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, CIT, MIL, ANC, NATIVITY, RELP, DEAR, DEYE, DREM, RAC1P, COW, ESR, WKHP, JWMNP, and PINCP. The sub-population over which the task is conducted is US residents with age between 18 and 35.
ACSEmployment
The goal of the ACSEmployment is to predict whether an individual is employed, given by the ESR column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, MIG, CIT, MIL, ANC, NATIVITY, RELP, DEAR, DEYE, DREM, and RAC1P. The sub-population over which the task is conducted is US residents with age between 16 and 90.
ACSTravelTime
The goal of the ACSTravelTime task is to predict whether a person’s commute time to work is greater than 20 minutes, given by the JWMNP column. The ACS columns used as features are: AGEP, SCHL, MAR, SEX, DIS, ESP, MIG, RELP, RAC1P, ST, CIT, OCCP, JWTR, and POVPIP. The sub-population over which the task is conducted is employed US residents with age greater than 16 years.
Supplementary Materials C folktexts package usage
The folktexts package is made available to the public via its open-source code repository1 and as a standalone package to be installed via the Python Package Index (PyPI). It is compatible with PyTorch models used locally, as well as with web-hosted models available through an API. The main user-facing classes are Benchmark, BenchmarkConfig, LLMClassifier, TaskMetadata, ColumnToText, and Dataset. The responsibilities of each class are ascribed as follows
  • The Benchmark class is responsible for running a benchmark task, which consists in obtaining risk scores from a given LLM on a given dataset, and evaluating those predictions on a variety of benchmark metrics.
  • The BenchmarkConfig class details all configurations of a benchmark (see Figure A10 for available options).
  • The LLMClassifier class is comprised of a transformers model, a tokenizer, and a task; and is responsible for producing risk scores given some tabular rows for the provided task.
  • The TaskMetadata class is responsible for defining a set of feature columns and target column, together with holding the corresponding column-to-text objects to map an entire tabular row to its natural-text representation. The benchmark ACS tasks instantiate a subclass named ACSTaskMetadata.
  • The ColumnToText class is responsible for producing meaningful natural-text representations of each possible value of a numeric or categorical column.
  • The Dataset class is responsible for holding tabular data and enabling reproducible manipulation of that data, such as splitting in train/test/validation, or filtering for a specified sub-population. The data used for the benchmark ACS tasks is provided by a subclass named ACSDataset.
Additionally, a command-line interface is provided to ease usability: The benchmark ACS tasks can be ran using the run_acs_benchmark executable. Figure A10 details each available flag. Further infromation and example notebooks can be found on github at: https://github.com/socialfoundations/folktexts.
usage:
run_acs_benchmark [-h] model MODEL results-dir RESULTS_DIR data-dir DATA_DIR [–task TASK] [–few-shot FEW_SHOT] [–batch-size BATCH_SIZE] [–context-size CONTEXT_SIZE] [–fit-threshold FIT_THRESHOLD] [–subsampling SUBSAMPLING] [–seed SEED] [–use-web-api-model] [–dont-correct-order-bias] [–numeric-risk-prompting] [–reuse-few-shot-examples] [–use-feature-subset USE_FEATURE_SUBSET]
[–use-population-filter USE_POPULATION_FILTER] [–logger-level {DEBUG,INFO,WARNING,ERROR,CRITICAL}]
Benchmark risk scores produced by a language model on ACS data.
options:
-h, help show this help message and exit
model MODEL [str] Model name or path to model saved on disk
results-dir RESULTS_DIR
[str] Directory under which this experiments results will be saved
data-dir DATA_DIR [str] Root folder to find datasets on
task TASK [str] Name of the ACS task to run the experiment on
few-shot FEW_SHOT [int] Use few-shot prompting with the given number of shots
batch-size BATCH_SIZE
[int] The batch size to use for inference
context-size CONTEXT_SIZE
[int] The maximum context size when prompting the LLM
fit-threshold FIT_THRESHOLD
[int] Whether to fit the prediction threshold, and on how many samples
subsampling SUBSAMPLING
[float] Which fraction of the dataset to use (if omitted will use all data)
seed SEED [int] Random seed to set for reproducibility
use-web-api-model [bool] Whether use a model hosted on a web API (instead of a local model)
dont-correct-order-bias
[bool] Whether to avoid correcting ordering bias, by default will correct it
numeric-risk-prompting
[bool] Whether to prompt for numeric risk-estimates instead of multiple-choice Q&A
reuse-few-shot-examples
[bool] Whether to reuse the same samples for few-shot prompting (or sample new ones every time)
use-feature-subset USE_FEATURE_SUBSET
[str] Optional subset of features to use for prediction, comma separated
use-population-filter USE_POPULATION_FILTER
[str] Optional population filter for this benchmark; must follow the format column_name=value to filter the dataset by a specific value.
logger-level {DEBUG,INFO,WARNING,ERROR,CRITICAL}
[str] The logging level to use for the experiment
Figure A10: Documentation for using folktexts package through the command-line interface. An executable named run_acs_benchmark is made available to run the standard ACS benchmark tasks with a variety of available customization options. Detailed documentation available at socialfoundations.github.io/folktexts/
Column Description Example
AGEP Age The individual’s age is: 42 years old.
COW Class of worker The individual’s current employment status is: Working for a non-profit organization.
SCHL Educational attainment The individual’s highest grade completed is: 12th grade.
MAR Marital status The individual’s marital status is: Married.
OCCP Occupation The individual’s occupation is: Human Resources Manager.
POBP Place of birth The individual’s place of birth is: New Zealand.
RELP Relationship The individual’s relationship to the reference survey respondent in the household is: Brother or sister.
WKHP Work-hours per week The individual’s usual number of hours worked per week is: 40 hours.
SEX Sex The individual’s sex is: Female.
RAC1P Race The individual’s race is: Black or African American.
PINCP Total yearly income The individual’s total yearly income is: \$75,000.
CIT Citizenship status The individual’s citizenship status is: Naturalized US citizen.
DIS Disability status The individual has a disability.
ESP Employment status of parents The individual is living with two parents: both parents in labor force.
MIG Mobility (lived here 1 year ago) The individual lived in the same house 1 year ago.
MIL Military service The individual was on active duty in the past, but not currently.
PUBCOV Public health coverage The individual is covered by public health insurance.
ANC Ancestry The individual has single ancestry.
NATIVITY Nativity The individual is foreign born.
DEAR Hearing The individual has hearing difficulty.
DEYE Vision The individual does not have vision difficulty.
DREM Cognition The individual does not have cognitive difficulties.
ESR Employment status #2 The individual is not in the labor force.
ST State The individual lives in California.
FER Parenthood (1 year) The individual gave birth to a child within the past 12 months.
JWMNP Commute time The individual takes 45 minutes travelling to work every day.
JWTR Means of transport The individual’s means of transport to work is a bicycle.
POVPIP Income-to-poverty ratio The individual’s income to poverty ratio is 150%.
Table A5: Description of all column-to-text mappings implemented for ACS features. The variable part of each example is shown in typeset grey font. Details on each possible categorical value for each feature are available in the ACS PUMS data dictionary.5